### v4

# Pentagon relation in quantum cluster scattering diagrams

Abstract.
We formulate the pentagon relation for quantum dilogarithm elements in the structure group of a quantum cluster scattering diagram (QCSD). As an application, we establish the nonpositivity of a certain class of nonskew-symmetric QCSDs. Also, we explicitly present various consistency relations and their positivity for QCSDs of rank 2 completely or up to some degree, most of which are new in the literature.
Key words and phrases:
cluster algebra, scattering diagram, quantum dilogarithm
2020 Mathematics Subject Classification:
Primary 13F60
1. Introduction
In [GHKK18] the scattering diagram method was employed to study cluster algebras [FZ02], where all information of a cluster algebra or a cluster pattern is encoded in the corresponding cluster scattering diagram (CSD, for short). It was clarified implicitly in [GHKK18] and more explicitly in [Nak23] that the dilogarithm elements $\Psi[n]$, where $n$ is its normal vector, and the pentagon relation among them play prominent roles in a CSD.
Meanwhile, it has been known that cluster algebras admit natural quantizations [BZ05, Tra11, FG09a, FG09b]. The relation to the quantum dilogarithm [FK94] was also studied in [FG09b, Kel11, KN11, Nag11]. Naturally, quantum cluster scattering diagrams (QCSD, for short), which are quantum analogues of CSDs, were introduced and studied in [Man21, DM21], where quantum dilogarithm elements again play a primary role. Here and below, we distinguish the quantum dilogarithm $\boldsymbol{\Psi}_{q}(x)$ and the quantum dilogarithm elements $\Psi_{a,b}[n]$ to avoid the confusion. The former is a formal power series of an indeterminate $x$ with a parameter $q$ due to [Sch53, FK94], while the latter are certain elements of the structure group $G$ of a QCSD. It is well known that the quantum dilogarithm satisfies the pentagon identity [FK94, CF99], which is a quantum analogue of the pentagon identity (more often called Abel’s identity or the five-term relation) of the dilogarithm [Lew81].
Having this development in mind, in this paper we formulate the pentagon relation among the quantum dilogarithm elements (Theorem 2.7), which is the counterpart of the one for the dilogarithm elements for a CSD in [Nak23]. The definition of QCSDs is just as parallel to the one for CSDs thanks to a more general formulation of scattering diagrams by [KS14]. However, a quantum dilogarithm element $\Psi_{a,b}[n]$ carries some additional rational parameters (the quantum data) $a$ and $b$ compared with its classical counterpart $\Psi[n]$, and this is the source of the complication and the richness in the quantum case. (The situation is somewhat similar to the fusion of representations of quantum affine algebras with the spectral parameter.)
As an application of the pentagon relation, together with the combinatorics of the quantum data, we establish the nonpositivity of a certain class of nonskew-symmetric QCSDs (Theorem 5.6 and Corollary 5.7), which extends the results of [LLRZ14, DM21, CFMM20]. Also, we explicitly present various consistency relations and their positivity for QCSDs of rank 2 completely (Section 3.2) or up to some degree (Section 5.4), most of which are new in the literature.
Acknowledgment. We thank Peigen Cao for useful comments and especially for pointing out the error in the proof of Lemma 4.2 in the early manuscript. This work is partly supported by JSPS Grant Numbers JP16H03922 and JP22H01114.
2. Quantum dilogarithm elements and pentagon relation
We introduce quantum dilogarithm elements and the pentagon relation among them.
2.1. Structure group $G$
Here we introduce the underlying structure group $G$ of a QCSD following [Man21, DM21, CFMM20, KS14]. See also [GHKK18, Nak23] for parallel notions for a CSD.
Let $\Gamma$ be a fixed data consisting of the following:
  • a lattice $N$ of rank $r$,
  • a skew-symmetric bilinear form {,}$\{\cdot,\cdot\}\colon N\times N\rightarrow\mathbb{Q}$,
  • a sublattice $N^{\circ}\subset N$ of rank $r$ such that $\{N^{\circ},N\}\subset\mathbb{Z}$,
  • positive integers $\delta_{1}$, …, $\delta_{r}$ such that there is a basis $(e_{1},\dots,e_{r})$ of $N$, where $(\delta_{1}e_{1},\dots,\delta_{r}e_{r})$ is a basis of $N^{\circ}$,
  • $M=\mathrm{Hom}(N,\mathbb{Z})$ and $M^{\circ}=\mathrm{Hom}(N^{\circ},\mathbb{Z})$.
Let $M_{\mathbb{R}}=M\otimes_{\mathbb{Z}}\mathbb{R}$. Let $\langle n,m\rangle$ denote the canonical paring either for $N^{\circ}\times M^{\circ}$ or for $N\times M_{\mathbb{R}}$. For $n\in N$, $n\neq 0$, let $n^{\perp}:=\{\xi\in M_{\mathbb{R}}\mid\langle n,\xi\rangle=0\}$.
Let $\mathfrak{s}=(e_{1},\dots,e_{r})$ be a seed for $\Gamma$, which is a basis of $N$ such that $(\delta_{1}e_{1},\dots,\delta_{r}e_{r})$ is a basis of $N^{\circ}$. The dual bases of $M$ and $M^{\circ}$ are given by $(e_{1}^{*},\dots,e_{r}^{*})$ and $(f_{1},\dots,f_{r}):=(e_{1}^{*}/\delta_{1},\dots,e_{r}^{*}/\delta_{r})$, respectively. The initial exchange matrix $B=(b_{ij})$ of the corresponding quantum cluster algebra is given by
(2.1) $b_{ij}=\{\delta_{i}e_{i},e_{j}\}.$
Let
(2.2) $N^{+}=\Biggl{\{}\sum_{i=1}^{r}a_{i}e_{i}\mid a_{i}\in\mathbb{Z}_{\geq 0},\,\sum_{i=1}^{r}a_{i}>0\Biggr{\}}$
be the positive vectors of $N$ with respect to $\mathfrak{s}$. Let $N^{+}_{\mathrm{pr}}$ denote the set of primitive elements in $N^{+}$. The degree function $\deg\colon N^{+}\rightarrow\mathbb{Z}_{>0}$ is defined by $\deg(\sum_{i=1}^{r}a_{i}e_{i}):=\sum_{i=1}^{r}a_{i}$.
For $n\in N^{+}$, let $\delta(n)$ be the smallest positive rational number such that $\delta(n)n\in N^{\circ}$, which is called the normalization factor of $n$. For example, $\delta(e_{i})=\delta_{i}$. We have $\delta(n)\in\mathbb{Z}_{>0}$ and $\delta(tn)=\delta(n)/t$ for any $n\in N^{+}_{\mathrm{pr}}$ and $t\in\mathbb{Z}_{>0}$. We set $\delta_{0}$ as the least common multiple of $\delta_{1}$, …, $\delta_{r}$. Then, $\delta_{0}n\in N^{\circ}$ for any $n\in N^{+}$. It implies that {,}$\{\cdot,\cdot\}\in(1/\delta_{0})\mathbb{Z}$ and that $\delta_{0}/\delta(n)\in\mathbb{Z}_{>0}$.
Let $q$ be an indeterminate, and let $\mathbb{Q}(q^{1/\delta_{0}})$ be the rational function field of $q^{1/\delta_{0}}$. For any $\alpha\in(1/\delta_{0})\mathbb{Z}$, let
(2.3) $[\alpha]_{q}:=\frac{q^{\alpha}-q^{-\alpha}}{q-q^{-1}}\in\mathbb{Q}(q^{1/\delta_{0}})$
be the $q$-number, which has the limit
(2.4) $\lim_{q\rightarrow 1}[\alpha]_{q}=\alpha.$
Let $\mathfrak{g}$ be the $N^{+}$-graded Lie algebra over $\mathbb{Q}(q^{1/\delta_{0}})$ defined by
(2.5) $\mathfrak{g}=\bigoplus_{n\in N^{+}}\mathbb{Q}(q^{1/\delta_{0}})X_{n},\quad[X_{n},X_{n^{\prime}}]:=[\{n,n^{\prime}\}]_{q}X_{n+n^{\prime}},$
where the Jacobi identity is easily verified. Let $\widehat{\mathfrak{g}}$ be the completion of $\mathfrak{g}$ with respect to $\deg$, and let $G=\exp(\widehat{\mathfrak{g}})$ be the exponential group of $\widehat{\mathfrak{g}}$ whose product is defined by the Baker-Campbell-Hausdorff formula (e.g., [Jac79, §V.5]). We call $G$ the structure group for the forthcoming scattering diagrams.
Let $(N^{+})^{>\ell}=\{n\in N^{+}\mid\deg(n)>\ell\}$. Let $G^{>\ell}$ be the normal subgroup of $G$ consisting elements $\exp(\sum_{n\in(N^{+})^{>\ell}}c_{n}X_{n})$, where the sum may be infinite, and let $G^{\leq\ell}:=G/G^{>\ell}$ be its quotient. For each $n\in N_{\mathrm{pr}}^{+}$, let $G_{n}^{\parallel}$ be the abelian subgroup of $G$ consisting of elements $\exp(\sum_{j=1}^{\infty}c_{j}X_{jn})$. For any $g=\exp(X)$ and $c\in\mathbb{Q}(q^{1/\delta_{0}})$, the power $g^{c}\in G$ is defined by $\exp(cX)$.
Remark 2.1.
So far, the only difference from the classical case is the Lie bracket in (2.5). It is a quantum analogue of the Lie bracket
(2.6) $[X_{n},X_{n^{\prime}}]:=\{n,n^{\prime}\}X_{n+n^{\prime}}.$
for a CSD. In what follows all expressions involving $q$ converge to their classical counterparts in the limit $q\rightarrow 1$.
2.2. $y$-representation
Let $y$ be a formal symbol. Let $\mathbb{Q}(q^{1/\delta_{0}})[y]_{q}$ be the noncommutative and associative algebra over $\mathbb{Q}(q^{1/\delta_{0}})$ with generators $y^{n}$ ($n\in N^{+}\sqcup\{0\})$ and the relations
(2.7) $y^{n}y^{n^{\prime}}=q^{\{n,n^{\prime}\}}y^{n+n^{\prime}}.$
Let $\mathcal{R}_{q}(y)$ be the completion of $\mathbb{Q}(q^{1/\delta_{0}})[y]_{q}$ with respect to $\deg$. In other words, any element of $\mathcal{R}_{q}(y)$ is expressed as an infinite sum
(2.8) $\sum_{n\in Q}c_{n}y^{n}\quad(c_{n}\in\mathbb{Q}(q^{1/\delta_{0}})).$
Then, we define the action of $\widehat{\mathfrak{g}}$ on $\mathcal{R}_{q}(y)$ by
(2.9) $X_{n}(y^{n^{\prime}}):=$ [$\ [\{n,n^{\prime}\}]_{q}y^{n^{\prime}+n}$
(2.10) $=$ $\ \frac{q^{2\{n,n^{\prime}\}}-1}{q-q^{-1}}y^{n^{\prime}}y^{n}.$
It is easy to check that this is indeed an action of $\widehat{\mathfrak{g}}$ and also a derivation. Thus, it induces the action of $G$ on $\mathcal{R}_{q}(y)$ defined by
(2.11) $(\exp X)(y^{n})=\sum_{j=0}^{\infty}\frac{1}{j!}X^{j}(y^{n})\quad(X\in\widehat{\mathfrak{g}}).$
[…]
Remark 2.2.
If we replace the relation (2.7) with the standard one $y^{n}y^{n^{\prime}}=y^{n+n^{\prime}}$, the action (2.9) is still an action of $\widehat{\mathfrak{g}}$, but it is not a derivation. This explains the necessity of the relation (2.7).
2.3. Quantum dilogarithm elements and pentagon relation
Let $\boldsymbol{\Psi}_{q}(x)$ be the quantum dilogarithm by [Sch53, FV93, FK94],
(2.12) $\boldsymbol{\Psi}_{q}(x):=$ $\ \exp\biggr{(}\sum_{j=1}^{\infty}\frac{(-1)^{j+1}}{j(q^{j}-q^{-j})}x^{j}\biggl{)}\quad\in\mathbb{Q}(q)[[x]]$
(2.13) $=$ $\ \prod_{j=0}^{\infty}(1+q^{2j+1}x)^{-1}.$
The equality holds, because the both expressions satisfy the property
(2.14) $\boldsymbol{\Psi}_{q}(q^{2}x)=(1+qx)\boldsymbol{\Psi}_{q}(x),\quad\boldsymbol{\Psi}_{q}(0)=1,$
which uniquely determines $\boldsymbol{\Psi}_{q}(x)$. See, for example, [Kir95, FG09a] for further information.
The following definition is similar to the first expression (2.12) of $\boldsymbol{\Psi}_{q}(x)$.
Definition 2.3 (Quantum dilogarithm element).
For any $n\in N^{+}$, $a\in(1/\delta_{0})\mathbb{Z}_{>0}$, and $b\in(1/\delta_{0})\mathbb{Z}$, we define
(2.15) $\Psi_{a,b}[n]:=\exp\Biggl{(}\sum_{j=1}^{\infty}\frac{(-1)^{j+1}}{j[ja]_{q}}q^{jb}X_{jn}\Biggr{)}\in G_{n_{0}}^{\parallel},$
where $n_{0}\in N_{\mathrm{pr}}^{+}$ is the one such that $n=jn_{0}$ for some $j\in\mathbb{Z}_{>0}$. We call it a quantum dilogarithm element. For simplicity, we also write
(2.16) $\Psi_{a}[n]:=\Psi_{a,0}[n].$
We call the parameters $a$ and $b$ the quantum data of $\Psi_{a,b}[n]$. They control the interval and the shift of $\Psi_{a,b}[n]$, respectively, from the standard one $\Psi_{1,0}[n]$.
Remark 2.4.
The above $\Psi_{a,b}[n]$ is a quantum analogue of the dilogarithm element in the classical case [Nak23, §III.1.4]
(2.17) $\Psi[n]=\exp\Biggl{(}\sum_{j=1}^{\infty}\frac{(-1)^{j+1}}{j^{2}}X_{jn}\Biggr{)}.$
We have
(2.18) $\lim_{q\rightarrow 1}\Psi_{a,b}[n]=\Psi[n]^{1/a}.$
Note that the interval $a$ appears in the RHS as the inverse.
Proposition 2.5.
The group $G$ is generated by $\Psi_{a,b}[n]^{c}$ ($n\in N^{+}$, $a\in(1/\delta_{0})\mathbb{Z}_{>0}$, $b\in(1/\delta_{0})\mathbb{Z}$, $c\in\mathbb{Q}(q^{1/\delta_{0}})$) admitting the infinite product.
Proof.
By (2.15), any element of $G_{n_{0}}^{\parallel}$ is expressed as an infinite product of $\Psi_{a,b}[jn_{0}]^{c}$’s. ∎
By (2.10), a dilogarithm element $\Psi_{a,b}[n]$ acts on $\mathcal{R}_{q}(y)$ under $\rho_{y}$ as
(2.19) $\begin{split}\Psi_{a,b}[n](y^{n^{\prime}})&=\exp\Biggl{(}\,\sum_{j=1}^{\infty}\frac{(-1)^{j+1}}{j[ja]_{q}}q^{jb}X_{jn}\Biggr{)}(y^{n^{\prime}})\\ &=y^{n^{\prime}}\exp\Biggl{(}\sum_{j=1}^{\infty}\frac{q^{2j\{n,n^{\prime}\}}-1}{q^{2ja}-1}\frac{(-1)^{j+1}}{j}q^{ja}q^{jb}y^{jn}\Biggr{)}.\end{split}$
We consider two special situations, where the expression (2.19) is simplified.
(i). Fix $n$ in (2.19). Let $a=1/s\delta(n)$, where $s\in\mathbb{Z}_{>0}$ is a divisor of $\delta_{0}/\delta(n)$. Then, $a\in(1/\delta_{0})\mathbb{Z}_{>0}$, and also $\alpha:=\{s\delta(n)n,n^{\prime}\}$ is an integer. Then, we have
(2.20) $\begin{split}\frac{q^{2j\{n,n^{\prime}\}}-1}{q^{2ja}-1}&=\frac{q^{2\alpha j/s\delta(n)}-1}{q^{2j/s\delta(n)}-1}=\begin{cases}\displaystyle\sum_{p=0}^{\alpha-1}q^{2jp/s\delta(n)}&\alpha>0,\\ 0&\alpha=0,\\ \displaystyle-\sum_{p=1}^{-\alpha}q^{-2jp/s\delta(n)}&\alpha<0.\end{cases}\end{split}$
Thus, we obtain
(2.21) $\begin{split}&\quad\ \Psi_{1/s\delta(n),b}[n](y^{n^{\prime}})=\begin{cases}\displaystyle y^{n^{\prime}}\prod_{p=1}^{\alpha}(1+q^{(2p-1)/s\delta(n)}q^{b}y^{n})&\alpha>0,\\ y^{n^{\prime}}&\alpha=0,\\ \displaystyle y^{n^{\prime}}\prod_{p=1}^{-\alpha}(1+q^{-(2p-1)/s\delta(n)}q^{b}y^{n})^{-1}&\alpha<0.\end{cases}\end{split}$
Observe that this is the automorphism part of the Fock-Goncharov decomposition of mutations of the quantum $y$-variables in [FG09a, §3.3].
(ii). Fix $n$ and $n^{\prime}$ in (2.19). Suppose that $\{n,n^{\prime}\}\neq 0$, and let $a=|\{n,n^{\prime}\}|$. Then, we have
(2.22) $\begin{split}\frac{q^{2j\{n,n^{\prime}\}}-1}{q^{2ja}-1}&=\begin{cases}1&\{n,n^{\prime}\}>0,\\ \displaystyle-q^{-2ja}&\{n,n^{\prime}\}<0.\end{cases}\end{split}$
Thus, we obtain
(2.23) $\begin{split}&\quad\ \Psi_{a,b}[n](y^{n^{\prime}})=\begin{cases}\displaystyle y^{n^{\prime}}(1+q^{a}q^{b}y^{n})&\{n,n^{\prime}\}>0,\\ \displaystyle y^{n^{\prime}}(1+q^{-a}q^{b}y^{n})^{-1}&\{n,n^{\prime}\}<0.\end{cases}\end{split}$
Remark 2.6.
The same formula in the RHS of (2.21) is also obtained by the adjoint action of the quantum dilogarithm $\boldsymbol{\Psi}_{q^{1/s\delta(n)}}(q^{b}y^{n})$ on $\mathcal{R}_{q}(y)$ [FG09a]. In this sense we have the correspondence between $\Psi_{1/s\delta(n),b}[n]$ and the quantum dilogarithm $\boldsymbol{\Psi}_{q^{1/s\delta(n)}}(q^{b}y^{n})$.
We have the following important application of the formula (2.23).
Theorem 2.7.
Let $n_{1},n_{2}\in N^{+}$. The following relations hold in $G$.
(a). If $\{n_{2},n_{1}\}=0$, we have, for any $a_{i}$ and $b_{i}$,
(2.24) $\Psi_{a_{2},b_{2}}[n_{2}]\Psi_{a_{1},b_{1}}[n_{1}]=\Psi_{a_{1},b_{1}}[n_{1}]\Psi_{a_{2},b_{2}}[n_{2}].$
(b). (Pentagon relation. Cf. [FK94, §2]). If $\{n_{2},n_{1}\}=c$ $(c\in(1/\delta_{0})\mathbb{Z}_{>0})$, we have, for any $b_{i}$,
(2.25) $\Psi_{c,b_{2}}[n_{2}]\Psi_{c,b_{1}}[n_{1}]=\Psi_{c,b_{1}}[n_{1}]\Psi_{c,b_{1}+b_{2}}[n_{1}+n_{2}]\Psi_{c,b_{2}}[n_{2}].$
Proof.
The relation (2.24) is clear by (2.5). Let us prove the relation (2.25). Since it only involves $\Psi_{a,b}[n]$’s for the rank 2 sublattice $N^{\prime}$ of $N$ generated by $n_{1}$ and $n_{2}$, we concentrate on the subgroup $G^{\prime}$ of $G$ corresponding to $N^{\prime}$. Accordingly, we consider the $y$-representation of $G^{\prime}$ acting on the subalgebra of $\mathcal{R}_{q}(y)$ generated by $y^{n_{1}}$ and $y^{n_{2}}$. By the assumption $\{n_{2},n_{1}\}\neq 0$, this is faithful. Thus, one can prove (2.25) with this representation.
First, we consider the action on $y^{n_{1}}$. Note that $\{n_{1}+n_{2},n_{1}\}=-\{n_{1}+n_{2},n_{2}\}=c>0$. Then, thanks to the formula (2.23), we have
(2.26) $\begin{split}(\Psi_{c,b_{2}}[n_{2}]\Psi_{c,b_{1}}[n_{1}])(y^{n_{1}})&=\Psi_{c,b_{2}}[n_{2}](y^{n_{1}})\\ &=y^{n_{1}}(1+q^{c+b_{2}}y^{n_{2}}),\end{split}$
and also
(2.27) $\begin{split}&\quad\ (\Psi_{c,b_{1}}[n_{1}]\Psi_{c,b_{1}+b_{2}}[n_{1}+n_{2}]\Psi_{c,b_{2}}[n_{2}])(y^{n_{1}})\\ &=(\Psi_{c,b_{1}}[n_{1}]\Psi_{c,b_{1}+b_{2}}[n_{1}+n_{2}])(y^{n_{1}}(1+q^{c+b_{2}}y^{n_{2}}))\\ &=\Psi_{c,b_{1}}[n_{1}](y^{n_{1}}(1+q^{c+b_{1}+b_{2}}y^{n_{1}+n_{2}})\\ &\qquad\times(1+q^{c+b_{2}}y^{n_{2}}(1+q^{-c+b_{1}+b_{2}}y^{n_{1}+n_{2}})^{-1}))\\ &=\Psi_{c,b_{1}}[n_{1}](y^{n_{1}}(1+q^{c+b_{2}}y^{n_{2}}+q^{c+b_{1}+b_{2}}y^{n_{1}+n_{2}}))\\ &=y^{n_{1}}(1+q^{c+b_{2}}y^{n_{2}}(1+q^{-c+b_{1}}y_{1})^{-1}+q^{c+b_{1}+b_{2}}y^{n_{1}+n_{2}}(1+q^{-c+b_{1}}y_{1})^{-1})\\ &=y^{n_{1}}(1+q^{-c+b_{1}}y_{1}+q^{c+b_{2}}y^{n_{2}}+q^{b_{1}+b_{2}}y^{n_{2}}y^{n_{1}})(1+q^{-c+b_{1}}y_{1})^{-1}\\ &=y^{n_{1}}(1+q^{c+b_{2}}y^{n_{2}}).\end{split}$
Thus, they coincide. Next, we consider the action on $y^{n_{2}}$. Similarly, we have
(2.28) $\begin{split}&\quad\ (\Psi_{c,b_{2}}[n_{2}]\Psi_{c,b_{1}}[n_{1}])(y^{n_{2}})\\ &=\Psi_{c,b_{2}}[n_{2}](y^{n_{2}}(1+q^{-c+b_{1}}y^{n_{1}})^{-1})\\ &=y^{n_{2}}(1+q^{-c+b_{1}}y^{n_{1}}(1+q^{c+b_{2}}y^{n_{2}}))^{-1}\\ &=y^{n_{2}}(1+q^{-c+b_{1}}y^{n_{1}}+q^{b_{1}+b_{2}}y^{n_{1}}y^{n_{2}})^{-1},\end{split}$
and also
(2.29) $\begin{split}&\quad\ (\Psi_{c,b_{1}}[n_{1}]\Psi_{c,b_{1}+b_{2}}[n_{1}+n_{2}]\Psi_{c,b_{2}}[n_{2}])(y^{n_{2}})\\ &=(\Psi_{c,b_{1}}[n_{1}]\Psi_{c,b_{1}+b_{2}}[n_{1}+n_{2}])(y^{n_{2}})\\ &=\Psi_{c,b_{1}}[n_{1}](y^{n_{2}}(1+q^{-c+b_{1}+b_{2}}y^{n_{1}+n_{2}})^{-1})\\ &=y^{n_{2}}(1+q^{-c+b_{1}}y^{n_{1}})^{-1}(1+q^{-c+b_{1}+b_{2}}y^{n_{1}+n_{2}}(1+q^{-c+b_{1}}y_{1})^{-1})^{-1}\\ &=y^{n_{2}}(1+q^{-c+b_{1}}y^{n_{1}}+q^{b_{1}+b_{2}}y^{n_{1}}y^{n_{2}})^{-1}.\end{split}$
Again, they coincide. ∎
Also, the following operations to decompose and unify quantum dilogarithm elements are important for our purpose.
Proposition 2.8.
Let $p$ be a positive integer. The following equalities hold, where we assume $a/p\in(1/\delta_{0})\mathbb{Z}_{>0}$ in the second equality:
[…]
Proof.
Two equalities (2.30) and (2.31) are equivalent, and they do not depend on the choice of the sign therein. We show (2.30) for the sign $+$ as follows:
(2.32) $\begin{split}\Psi_{a,b}[n]&=\exp\Biggl{(}\sum_{j=1}^{\infty}\frac{(-1)^{j+1}}{j[ja]_{q}}\frac{q^{jpa}-q^{-jpa}}{q^{jpa}-q^{-jpa}}q^{bj}X_{jn}\Biggr{)}\\ &=\exp\Biggl{(}\sum_{j=1}^{\infty}\frac{(-1)^{j+1}}{j[jpa]_{q}}\frac{q^{jpa}-q^{-jpa}}{q^{ja}-q^{-ja}}q^{bj}X_{jn}\Biggr{)}\\ &=\exp\Biggl{(}\sum_{t=1}^{p}\sum_{j=1}^{\infty}\frac{(-1)^{j+1}}{j[jpa]_{q}}q^{j(2t-p-1)a}q^{bj}X_{jn}\Biggr{)}\\ &=\prod_{t=1}^{p}\Psi_{pa,b+(2t-p-1)a}[n].\end{split}$
3. Quantum cluster scattering diagrams
We introduce quantum cluster scattering diagrams. Examples in rank 2 are also given.
3.1. Quantum cluster scattering diagrams
The following definitions for (quantum) scattering diagrams are just the same as the classical one in [GHKK18], by replacing the structure group $G$ with the one in Section 2.1.
A wall $\mathbf{w}=(\mathfrak{d},g)_{n}$ for $\mathfrak{s}$ is a triplet with $n\in N_{\mathrm{pr}}^{+}$, a cone $\mathfrak{d}\subset n^{\perp}\subset M_{\mathbb{R}}$ of codimension 1, and $g\in G_{n}^{\parallel}$. We call $n$, $\mathfrak{d}$, $g$, the normal vector, the support, the wall element of $\mathbf{w}$, respectively. Let $p^{*}\colon N\rightarrow M^{\circ}\subset M_{\mathbb{R}}$, $n\mapsto\{\cdot,n\}$. We say that a wall $\mathbf{w}=(\mathfrak{d},g)_{n}$ is incoming if $p^{*}(n)\in\mathfrak{d}$ holds.
Definition 3.1 (Scattering diagram).
A scattering diagram $\mathfrak{D}=\{\mathbf{w}_{\lambda}=(\mathfrak{d}_{\lambda},g_{{\lambda}})_{n_{\lambda}}\}_{\lambda\in\Lambda}$ for $\mathfrak{s}$ is a collection of walls for $\mathfrak{s}$ satisfying the following finiteness condition: For any degree $\ell$, there are only finitely many walls such that $\pi_{\ell}(g_{\lambda})\neq\mathrm{id}$, where $\pi_{\ell}\colon G\rightarrow G^{\leq\ell}$ is the canonical projection.
For a scattering diagram $\mathfrak{D}$, we define
(3.1) $\mathrm{Supp}(\mathfrak{D})$ $=\bigcup_{\lambda\in\Lambda}\mathfrak{d}_{\lambda},\quad\mathrm{Sing}(\mathfrak{D})=\bigcup_{\lambda\in\Lambda}\partial\mathfrak{d}_{\lambda}\cup\bigcup_{\scriptstyle\lambda,\lambda^{\prime}\in\Lambda\atop\scriptstyle\dim\mathfrak{d}_{\lambda}\cap\mathfrak{d}_{\lambda^{\prime}}=r-2}\mathfrak{d}_{\lambda}\cap\mathfrak{d}_{\lambda^{\prime}}.$
A curve $\gamma\colon[0,1]\rightarrow M_{\mathbb{R}}$ is admissible for $\mathfrak{D}$ if it satisfies the following properties:
  • (1)
    The end points of $\gamma$ are in $M_{\mathbb{R}}\setminus\mathrm{Supp}(\mathfrak{D})$.
  • (2)
    It is a smooth curve, and it intersects $\mathrm{Supp}(\mathfrak{D})$ transversally.
  • (3)
    $\gamma$ does not intersect $\mathrm{Sing}(\mathfrak{D})$.
For any admissible curve $\gamma$, the path-ordered product $\mathfrak{p}_{\gamma,\mathfrak{D}}\in G$ is defined as the product of the wall elements $g_{{\lambda}}^{\epsilon_{\lambda}}$ of walls $\mathbf{w}_{\lambda}$ of $\mathfrak{D}$ intersected by $\gamma$ in the order of intersection, where $\epsilon_{\lambda}$ is the intersection sign defined by
(3.2) $\epsilon_{\lambda}=\begin{cases}1&\langle n_{\lambda},\gamma^{\prime}\rangle<0,\\ -1&\langle n_{\lambda},\gamma^{\prime}\rangle>0,\end{cases}$
and $\gamma^{\prime}$ is the velocity vector of $\gamma$ at the wall $\mathbf{w}_{\lambda}$. The product $\mathfrak{p}_{\gamma,\mathfrak{D}}$ is an infinite one in general, and it is well-defined in $G$ due to the finiteness condition. We say that a pair of scattering diagrams $\mathfrak{D}$ and $\mathfrak{D}^{\prime}$ are equivalent if $\mathfrak{p}_{\gamma,\mathfrak{D}}=\mathfrak{p}_{\gamma,\mathfrak{D}^{\prime}}$ for any admissible curve $\gamma$ for both $\mathfrak{D}$ and $\mathfrak{D}^{\prime}$. We say that a scattering diagram $\mathfrak{D}$ is consistent if $\mathfrak{p}_{\gamma,\mathfrak{D}}=\mathrm{id}$ for any admissible loop (i.e., closed curve) $\gamma$ for $\mathfrak{D}$.
The following definition is due to [Man21].
Definition 3.2 (Quantum cluster scattering diagram).
A quantum cluster scattering diagram $\mathfrak{D}^{q}_{\mathfrak{s}}$ (QCSD, for short) for $\mathfrak{s}$ is a consistent scattering diagram whose set of incoming walls are given by
(3.3) $\mathrm{In}_{\mathfrak{s}}:=\{(e_{i}^{\perp},\Psi_{1/\delta_{i}}[e_{i}])_{e_{i}}\mid i=1,\dots,r\}.$
Theorem 3.3 ([GHKK18, Theorems 1.12], [KS14, Theorem 2.1.6], [DM21, Theorem 2.13]).
There exists a QCSD $\mathfrak{D}^{q}_{\mathfrak{s}}$ uniquely up to equivalence.
Proof.
The proof of [GHKK18, Theorem 1.12], which originates in [KS14, §2], is also applicable to this case. See also [Nak23, §III.3], which is closer to the present setting. ∎
Thanks to Proposition 2.5, we may assume that, up to equivalence, wall elements of $\mathfrak{D}_{\mathfrak{s}}^{q}$ are given by $\mathbb{Q}(q^{1/\delta_{0}})$-powers of quantum dilogarithm elements $\Psi_{a,b}[n]^{c}$ ($n\in N^{+}$, $c\in\mathbb{Q}(q^{1/\delta_{0}})$).
Remark 3.4.
To each admissible loop $\gamma$, the consistency condition $\mathfrak{p}_{\gamma,\mathfrak{D}}=\mathrm{id}$ gives a relation among quantum dilogarithm elements. Under the correspondence in Remark 2.6 this relation is translated as an identity for the quantum dilogarithm (quantum dilogarithm identity), which is a generalization of a quantum dilogarithm identity associated with a period of a quantum cluster pattern studied in [FG09b, Kel11, KN11, Nag11]. It is also the quantum counterpart of the (classical) dilogarithm identity for the Rogers dilogarithm in [Nak21].
3.2. Rank 2 examples
Let us demonstrate to construct the rank 2 QCSDs of finite and affine types based on the pentagon relation (2.25) in the spirit of [Nak23, §III.2.2, §III.3.5] for their classical counterparts.
3.2.1. Finite type
Here we consider the case where {,}$\{\cdot,\cdot\}\neq 0$. As in the classical case [Nak23, §III.1.5], without loosing generality, we may assume that $\{e_{2},e_{1}\}=1$ and $\delta_{2}\geq\delta_{1}$. Thus, the initial exchange matrix $B$ in (2.1) is given by
(3.6) $B=\begin{pmatrix}0&-\delta_{1}\\ \delta_{2}&0\end{pmatrix}.$
Let $e_{1}^{*},e_{2}^{*}\in M$ be the dual basis of $e_{1},e_{2}\in N$. Accordingly, let $f_{1}=e_{1}^{*}/\delta_{1},f_{2}=e_{2}^{*}/\delta_{2}\in M^{\circ}$ be the dual basis of $\delta_{1}e_{1},\delta_{2}e_{2}\in N^{\circ}$. Let $\mathbf{e}_{1},\mathbf{e}_{2}\in\mathbb{Z}^{2}$ be the unit vectors. We identify $N\simeq\mathbb{Z}^{2}$, $e_{i}\mapsto\mathbf{e}_{i}$, and $M_{\mathbb{R}}\simeq\mathbb{R}^{2}$, $f_{i}\mapsto\mathbf{e}_{i}$. Then, we have $\{\mathbf{n}^{\prime},\mathbf{n}\}=n^{\prime}_{2}n_{1}-n^{\prime}_{1}n_{2}$. Also, the canonical paring $\langle n,z\rangle\colon N\times M_{\mathbb{R}}\rightarrow\mathbb{R}$ is given by the corresponding vectors $\mathbf{n}$ and $\mathbf{z}$ as
(3.9) $\langle\mathbf{n},\mathbf{z}\rangle=\mathbf{n}^{T}\begin{pmatrix}\delta_{1}^{-1}&0\\ 0&\delta_{2}^{-1}\end{pmatrix}\mathbf{z}.$
For $\mathbf{n}=(n_{1},n_{2})$, we write
(3.14) $\begin{bmatrix}n_{1}\\ n_{2}\end{bmatrix}_{a,b}:=\Psi_{a,b}[\mathbf{n}],\quad\begin{bmatrix}n_{1}\\ n_{2}\end{bmatrix}_{a}:=\Psi_{a}[\mathbf{n}].$
[…]
Figure 1. Rank 2 QCSDs of finite type.
(a). Type $A_{2}$. Let $(\delta_{1},\delta_{2})=(1,1)$. Since $\{\mathbf{e}_{2},\mathbf{e}_{1}\}=1$, we apply the pentagon relation (2.25) with $c=1$, and we have
(3.25) $\begin{bmatrix}0\\ 1\end{bmatrix}_{1,0}\begin{bmatrix}1\\ 0\end{bmatrix}_{1,0}=\begin{bmatrix}1\\ 0\end{bmatrix}_{1,0}\begin{bmatrix}1\\ 1\end{bmatrix}_{1,0}\begin{bmatrix}0\\ 1\end{bmatrix}_{1,0}.$
This equality is naturally interpreted as a consistency relation of the QCSD of rank 2 in Figure 1 (a). Namely, it consists of three walls
(3.26) $(\mathbf{e}_{1}^{\perp},\Psi_{1}[\mathbf{e}_{1}])_{\mathbf{e}_{1}},\quad(\mathbf{e}_{2}^{\perp},\Psi_{1}[\mathbf{e}_{2}])_{\mathbf{e}_{2}},\quad(\mathbb{R}_{\geq 0}(1,-1),\Psi_{1}[(1,1)])_{(1,1)}.$
The LHS of the equality (3.25) is the path-ordered product $\mathfrak{p}_{\gamma_{1},\mathfrak{D}}$ along $\gamma_{1}$, while the RHS is the one along $\gamma_{2}$ in Figure 1.
(b). Type $B_{2}$. Let $(\delta_{1},\delta_{2})=(1,2)$. To use the pentagon relation (2.25) with $c=1$, we first apply the fission (2.30) for $\Psi_{1/2}(\mathbf{e}_{2})$ with $p=2$. Then apply the pentagon relation repeatedly for adjacent pairs, and apply the fusion (2.31) in the end. We have
(3.27) $\begin{split}\begin{bmatrix}0\\ 1\end{bmatrix}_{\frac{1}{2},0}\begin{bmatrix}1\\ 0\end{bmatrix}_{1,0}&=\begin{bmatrix}0\\ 1\end{bmatrix}_{1,-\frac{1}{2}}\begin{bmatrix}0\\ 1\end{bmatrix}_{1,\frac{1}{2}}\begin{bmatrix}1\\ 0\end{bmatrix}_{1,0}=\begin{bmatrix}0\\ 1\end{bmatrix}_{1,-\frac{1}{2}}\begin{bmatrix}1\\ 0\end{bmatrix}_{1,0}\begin{bmatrix}1\\ 1\end{bmatrix}_{1,\frac{1}{2}}\begin{bmatrix}0\\ 1\end{bmatrix}_{1,\frac{1}{2}}\\ &=\begin{bmatrix}1\\ 0\end{bmatrix}_{1,0}\begin{bmatrix}1\\ 1\end{bmatrix}_{1,-\frac{1}{2}}\begin{bmatrix}0\\ 1\end{bmatrix}_{1,-\frac{1}{2}}\begin{bmatrix}1\\ 1\end{bmatrix}_{1,\frac{1}{2}}\begin{bmatrix}0\\ 1\end{bmatrix}_{1,\frac{1}{2}}\\ &=\begin{bmatrix}1\\ 0\end{bmatrix}_{1,0}\begin{bmatrix}1\\ 1\end{bmatrix}_{1,-\frac{1}{2}}\begin{bmatrix}1\\ 1\end{bmatrix}_{1,\frac{1}{2}}\begin{bmatrix}1\\ 2\end{bmatrix}_{1,0}\begin{bmatrix}0\\ 1\end{bmatrix}_{1,-\frac{1}{2}}\begin{bmatrix}0\\ 1\end{bmatrix}_{1,\frac{1}{2}}\\ &=\begin{bmatrix}1\\ 0\end{bmatrix}_{1,0}\begin{bmatrix}1\\ 1\end{bmatrix}_{\frac{1}{2},0}\begin{bmatrix}1\\ 2\end{bmatrix}_{1,0}\begin{bmatrix}0\\ 1\end{bmatrix}_{\frac{1}{2},0}.\end{split}$
This equality is naturally interpreted as a consistency relation of the QCSD in Figure 1 (b), which consists of four walls
(3.28) $(\mathbf{e}_{1}^{\perp},\Psi_{1}[\mathbf{e}_{1}])_{\mathbf{e}_{1}},\quad(\mathbf{e}_{2}^{\perp},\Psi_{1/2}[\mathbf{e}_{2}])_{\mathbf{e}_{2}},$
(3.29) $(\mathbb{R}_{\geq 0}(1,-2),\Psi_{1/2}[(1,1)])_{(1,1)},\quad(\mathbb{R}_{\geq 0}(1,-1),\Psi_{1}[(1,2)])_{(1,2)}.$
(c). Type $G_{2}$. Let $(\delta_{1},\delta_{2})=(1,3)$. In the same way, we obtain
(3.30) $\begin{split}&\quad\,\begin{bmatrix}0\\ 1\end{bmatrix}_{\frac{1}{3},0}\begin{bmatrix}1\\ 0\end{bmatrix}_{1,0}=\begin{bmatrix}0\\ 1\end{bmatrix}_{1,-\frac{2}{3}}\begin{bmatrix}0\\ 1\end{bmatrix}_{1,0}\begin{bmatrix}0\\ 1\end{bmatrix}_{1,\frac{2}{3}}\begin{bmatrix}1\\ 0\end{bmatrix}_{1,0}\\ &=\begin{bmatrix}1\\ 0\end{bmatrix}_{1,0}\begin{bmatrix}1\\ 1\end{bmatrix}_{1,-\frac{2}{3}}\begin{bmatrix}1\\ 1\end{bmatrix}_{1,0}\begin{bmatrix}1\\ 1\end{bmatrix}_{1,\frac{2}{3}}\begin{bmatrix}2\\ 3\end{bmatrix}_{1,0}\\ &\qquad\times\begin{bmatrix}1\\ 2\end{bmatrix}_{1,-\frac{2}{3}}\begin{bmatrix}1\\ 2\end{bmatrix}_{1,0}\begin{bmatrix}1\\ 2\end{bmatrix}_{1,\frac{2}{3}}\begin{bmatrix}1\\ 3\end{bmatrix}_{1,0}\begin{bmatrix}0\\ 1\end{bmatrix}_{1,-\frac{2}{3}}\begin{bmatrix}0\\ 1\end{bmatrix}_{1,0}\begin{bmatrix}0\\ 1\end{bmatrix}_{1,\frac{2}{3}}\\ &=\begin{bmatrix}1\\ 0\end{bmatrix}_{1,0}\begin{bmatrix}1\\ 1\end{bmatrix}_{\frac{1}{3},0}\begin{bmatrix}2\\ 3\end{bmatrix}_{1,0}\begin{bmatrix}1\\ 2\end{bmatrix}_{\frac{1}{3},0}\begin{bmatrix}1\\ 3\end{bmatrix}_{1,0}\begin{bmatrix}0\\ 1\end{bmatrix}_{\frac{1}{3},0}.\end{split}$
This equality is naturally interpreted as a consistency relation of the QCSD in Figure 1 (c), which consists of six walls
(3.31) $(\mathbf{e}_{1}^{\perp},\Psi_{1}[\mathbf{e}_{1}])_{\mathbf{e}_{1}},\quad(\mathbf{e}_{2}^{\perp},\Psi_{1/3}[\mathbf{e}_{2}])_{\mathbf{e}_{2}},$
(3.32) $(\mathbb{R}_{\geq 0}(1,-3),\Psi_{{1/3}}[(1,1)])_{(1,1)},\quad(\mathbb{R}_{\geq 0}(1,-2),\Psi_{1}[(2,3)])_{(2,3)},$
(3.33) $(\mathbb{R}_{\geq 0}(2,-3),\Psi_{{1/3}}[(1,2)])_{(1,2)},\quad(\mathbb{R}_{\geq 0}(1,-1),\Psi_{1}[(1,3)])_{(1,3)}.$
Note that, in the RHSs of (3.25), (3.27), and (3.30), all factors have the form $\Psi_{1/\delta(n)}[n]$.
3.2.2. Affine type
Here we consider the affine type, where $\delta_{1}\delta_{2}=4$ for $B$ in (2.1). There are two cases.
[…]
Figure 2. Rank 2 QCSDs of affine type.
(a). Type $A_{1}^{(1)}$. Let $(\delta_{1},\delta_{2})=(2,2)$. Note that, for any primitive $n\in N_{\mathrm{pr}}^{+}$, we have $\delta(n)=2$. The following description of the CSD $\mathfrak{D}_{\mathfrak{s}}$ is known [GHKK18, Rei09, Rea20, Nak23, Mat21]: The walls of $\mathfrak{D}_{\mathfrak{s}}$ are given by
(3.34) $(\mathbf{e}_{1}^{\perp},\Psi[\mathbf{e}_{1}]^{2})_{\mathbf{e}_{1}},\quad(\mathbf{e}_{2}^{\perp},\Psi[\mathbf{e}_{2}]^{2})_{\mathbf{e}_{2}},$
(3.35) $(\mathbb{R}_{\geq 0}(p,-p-1),\Psi[(p+1,p)]^{2})_{(p+1,p)}\quad(p\in\mathbb{Z}_{>0}),$
(3.36) $(\mathbb{R}_{\geq 0}(p+1,-p),\Psi[(p,p+1)]^{2})_{(p,p+1)}\quad(p\in\mathbb{Z}_{>0}),$
(3.37) $(\mathbb{R}_{\geq 0}(1,-1),\Psi[2^{j}\mathbf{n}_{0}]^{2^{2-j}})_{\mathbf{n}_{0}}\quad(j\in\mathbb{Z}_{\geq 0}),$
where $\mathbf{n}_{0}=(1,1)$. Note that $\delta((p,p+1))=2$ and $\delta(2^{j}\mathbf{n}_{0})=2^{1-j}$. See Figure 2 (a).
Again, we apply the fission and the pentagon relation to the element $\Psi_{1/2}[\mathbf{e}_{2}]\Psi_{1/2}[\mathbf{e}_{1}]$ and obtain the equality
(3.38) $\begin{split}&\quad\,\begin{bmatrix}0\\ 1\end{bmatrix}_{\frac{1}{2},0}\begin{bmatrix}1\\ 0\end{bmatrix}_{\frac{1}{2},0}=\begin{bmatrix}0\\ 1\end{bmatrix}_{1,-\frac{1}{2}}\begin{bmatrix}0\\ 1\end{bmatrix}_{1,\frac{1}{2}}\begin{bmatrix}1\\ 0\end{bmatrix}_{1,-\frac{1}{2}}\begin{bmatrix}1\\ 0\end{bmatrix}_{1,\frac{1}{2}}\\ &=\begin{bmatrix}1\\ 0\end{bmatrix}_{1,-\frac{1}{2}}\begin{bmatrix}1\\ 1\end{bmatrix}_{1,-1}\begin{bmatrix}1\\ 1\end{bmatrix}_{1,0}\begin{bmatrix}1\\ 2\end{bmatrix}_{1,-\frac{1}{2}}\begin{bmatrix}1\\ 0\end{bmatrix}_{1,\frac{1}{2}}\begin{bmatrix}1\\ 1\end{bmatrix}_{1,0}\begin{bmatrix}1\\ 1\end{bmatrix}_{1,1}\begin{bmatrix}1\\ 2\end{bmatrix}_{1,\frac{1}{2}}\begin{bmatrix}0\\ 1\end{bmatrix}_{1,-\frac{1}{2}}\begin{bmatrix}0\\ 1\end{bmatrix}_{1,\frac{1}{2}}.\end{split}$
At this moment, the expression in the RHS is not yet ordered to be presented by a scattering diagram. Here, we say that the above product is ordered (resp. anti-ordered) if, for any adjacent pair $[\mathbf{n}^{\prime}]_{a^{\prime},b^{\prime}}[\mathbf{n}]_{a,b}$, $\{\mathbf{n}^{\prime},\mathbf{n}\}=n^{\prime}_{2}n_{1}-n^{\prime}_{1}n_{2}\leq 0$ (resp. $\{\mathbf{n}^{\prime},\mathbf{n}\}\geq 0$) holds. Equivalently, if we view $[\mathbf{n}]$ as a fraction $n_{1}/n_{2}$, then, the numbers should be aligned in the decreasing order (resp. increasing order) form left to right. The LHS of (3.38) is anti-ordered. To make the RHS of (3.38) ordered, we need to interchange the factors for $\mathbf{n}^{\prime}=(1,2)$ and $\mathbf{n}=(1,0)$ in the middle, where $\{\mathbf{n}^{\prime},\mathbf{n}\}=2$. As the lowest approximation, we consider modulo $G^{>3}$. Then, $[1,2]_{1,-1/2}$ commutes with other factors, and we have, modulo $G^{>3}$,
(3.55) $\begin{bmatrix}0\\ 1\end{bmatrix}_{\frac{1}{2},0}\begin{bmatrix}1\\ 0\end{bmatrix}_{\frac{1}{2},0}$ $\equiv\begin{bmatrix}1\\ 0\end{bmatrix}_{\frac{1}{2},0}\begin{bmatrix}2\\ 1\end{bmatrix}_{\frac{1}{2},0}\begin{bmatrix}1\\ 1\end{bmatrix}_{\frac{1}{2},-\frac{1}{2}}\begin{bmatrix}1\\ 1\end{bmatrix}_{\frac{1}{2},\frac{1}{2}}\begin{bmatrix}1\\ 2\end{bmatrix}_{\frac{1}{2},0}\begin{bmatrix}0\\ 1\end{bmatrix}_{\frac{1}{2},0}.$
To proceed to higher degree, we now apply the pentagon relation (2.25) with $c=2$ to the pair $[1,2]_{1,-1/2}$ and $[1,0]_{1,1/2}$ in (3.38). Then, we have
(3.56) $\begin{split}&\quad\,\begin{bmatrix}1\\ 2\end{bmatrix}_{1,-\frac{1}{2}}\begin{bmatrix}1\\ 0\end{bmatrix}_{1,\frac{1}{2}}=\begin{bmatrix}1\\ 2\end{bmatrix}_{2,-\frac{3}{2}}\begin{bmatrix}1\\ 2\end{bmatrix}_{2,\frac{1}{2}}\begin{bmatrix}1\\ 0\end{bmatrix}_{2,-\frac{1}{2}}\begin{bmatrix}1\\ 0\end{bmatrix}_{2,\frac{3}{2}}\\ &=\begin{bmatrix}1\\ 0\end{bmatrix}_{2,-\frac{1}{2}}\begin{bmatrix}2\\ 2\end{bmatrix}_{2,-2}\begin{bmatrix}2\\ 2\end{bmatrix}_{2,0}\begin{bmatrix}3\\ 4\end{bmatrix}_{2,-\frac{3}{2}}\begin{bmatrix}1\\ 0\end{bmatrix}_{2,\frac{3}{2}}\begin{bmatrix}2\\ 2\end{bmatrix}_{2,0}\begin{bmatrix}2\\ 2\end{bmatrix}_{2,2}\begin{bmatrix}3\\ 4\end{bmatrix}_{2,\frac{1}{2}}\begin{bmatrix}1\\ 2\end{bmatrix}_{2,-\frac{3}{2}}\begin{bmatrix}1\\ 2\end{bmatrix}_{2,\frac{1}{2}}.\end{split}$
This is parallel to (3.38) but with different quantum data. As the next approximation, we consider modulo $G^{>7}$. Then, $[3,4]_{2,-3/2}$ commutes with other factors, and we have, modulo $G^{>7}$,
(3.73) $\begin{bmatrix}1\\ 2\end{bmatrix}_{1,-\frac{1}{2}}\begin{bmatrix}1\\ 0\end{bmatrix}_{1,\frac{1}{2}}$ $\equiv\begin{bmatrix}1\\ 0\end{bmatrix}_{1,\frac{1}{2}}\begin{bmatrix}3\\ 2\end{bmatrix}_{1,\frac{1}{2}}\begin{bmatrix}2\\ 2\end{bmatrix}_{1,-1}\begin{bmatrix}2\\ 2\end{bmatrix}_{1,1}\begin{bmatrix}3\\ 4\end{bmatrix}_{1,-\frac{1}{2}}\begin{bmatrix}1\\ 2\end{bmatrix}_{1,-\frac{1}{2}}.$
Then, we plug it into (3.38), and apply the pentagon relation, and we have, modulo $G^{>7}$,
(3.74) $\begin{split}\begin{bmatrix}0\\ 1\end{bmatrix}_{\frac{1}{2},0}\begin{bmatrix}1\\ 0\end{bmatrix}_{\frac{1}{2},0}&\equiv\begin{bmatrix}1\\ 0\end{bmatrix}_{\frac{1}{2},0}\begin{bmatrix}2\\ 1\end{bmatrix}_{\frac{1}{2},0}\begin{bmatrix}3\\ 2\end{bmatrix}_{\frac{1}{2},0}\begin{bmatrix}4\\ 3\end{bmatrix}_{\frac{1}{2},0}\begin{bmatrix}1\\ 1\end{bmatrix}_{\frac{1}{2},-\frac{1}{2}}\begin{bmatrix}1\\ 1\end{bmatrix}_{\frac{1}{2},\frac{1}{2}}\\ &\quad\ \times\begin{bmatrix}2\\ 2\end{bmatrix}_{1,-1}\begin{bmatrix}2\\ 2\end{bmatrix}_{1,1}\begin{bmatrix}3\\ 4\end{bmatrix}_{\frac{1}{2},0}\begin{bmatrix}2\\ 3\end{bmatrix}_{\frac{1}{2},0}\begin{bmatrix}1\\ 2\end{bmatrix}_{\frac{1}{2},0}\begin{bmatrix}0\\ 1\end{bmatrix}_{\frac{1}{2},0}.\end{split}$
By continuing the procedure modulo $G^{>2^{\ell}-1}$, one can naturally guess that the relation converges to the following one:
(3.75) $\begin{split}\begin{bmatrix}0\\ 1\end{bmatrix}_{\frac{1}{2},0}\begin{bmatrix}1\\ 0\end{bmatrix}_{\frac{1}{2},0}&=\begin{bmatrix}1\\ 0\end{bmatrix}_{\frac{1}{2},0}\begin{bmatrix}2\\ 1\end{bmatrix}_{\frac{1}{2},0}\begin{bmatrix}3\\ 2\end{bmatrix}_{\frac{1}{2},0}\cdots\Biggl{(}\prod_{j=0}^{\infty}\begin{bmatrix}2^{j}\\ 2^{j}\end{bmatrix}_{2^{j-1},-2^{j-1}}\begin{bmatrix}2^{j}\\ 2^{j}\end{bmatrix}_{2^{j-1},2^{j-1}}\Biggr{)}\\ &\qquad\times\cdots\begin{bmatrix}2\\ 3\end{bmatrix}_{\frac{1}{2},0}\begin{bmatrix}1\\ 2\end{bmatrix}_{\frac{1}{2},0}\begin{bmatrix}0\\ 1\end{bmatrix}_{\frac{1}{2},0}.\end{split}$
Indeed, one can prove the relation (3.75) completely based on the pentagon relation by modifying the proof of [Mat21] for the classical case. The detail will be found in Appendix A. This is the simplest example of consistency relations involving infinite product. The formula appeared in [DGS11, Eq. (1.3)] without proof in an alternative setting of quantum dilogarithms. See also an alternative derivation in view of the quantum affine algebra of type $A_{1}^{(1)}$ [Sug22, Theorem 5.1].
(b). Type $A_{2}^{(2)}$. Let $(\delta_{1},\delta_{2})=(1,4)$. The situation is parallel to $A_{1}^{(1)}$ though a little more complicated. The following description of the CSD $\mathfrak{D}_{\mathfrak{s}}$ is known [Rea20, Nak23, Mat21]: The walls of $\mathfrak{D}_{\mathfrak{s}}$ are given by
(3.76) $(\mathbf{e}_{1}^{\perp},\Psi[\mathbf{e}_{1}])_{\mathbf{e}_{1}},\quad(\mathbf{e}_{2}^{\perp},\Psi[\mathbf{e}_{2}]^{4})_{\mathbf{e}_{2}},$
(3.77) $(\mathbb{R}_{\geq 0}(p,-2p-1),\Psi[(2p+1,4p)])_{(2p+1,4p)}\quad(p\in\mathbb{Z}_{>0}),$
(3.78) $(\mathbb{R}_{\geq 0}(2p-1,-4p),\Psi[(p,2p-1)]^{4})_{(p,2p-1)}\quad(p\in\mathbb{Z}_{>0}),$
(3.79) $(\mathbb{R}_{\geq 0}(p,-2p+1),\Psi[(2p-1,4p)])_{(2p-1,4p)}\quad(p\in\mathbb{Z}_{>0}),$
(3.80) $(\mathbb{R}_{\geq 0}(2p+1,-4p),\Psi[(p,2p+1)]^{4})_{(p,2p+1)}\quad(p\in\mathbb{Z}_{>0}),$
(3.81) $(\mathbb{R}_{\geq 0}(1,-2),\Psi[\mathbf{n}_{0}]^{6})_{\mathbf{n}_{0}},\quad(\mathbb{R}_{\geq 0}(1,-2),\Psi[2^{j}\mathbf{n}_{0}]^{2^{2-j}})_{\mathbf{n}_{0}}\quad(j\in\mathbb{Z}_{>0}),$
where $\mathbf{n}_{0}=(1,2)$. Note that $\delta(2^{j}\mathbf{n}_{0})=2^{1-j}$. See Figure 2 (b).
Again, we apply the fission and the pentagon relation to the element $\Psi_{1/4}[\mathbf{e}_{2}]\Psi_{1}[\mathbf{e}_{1}]$ and obtain the equality
(3.82) $\begin{split}&\quad\,\begin{bmatrix}0\\ 1\end{bmatrix}_{\frac{1}{4},0}\begin{bmatrix}1\\ 0\end{bmatrix}_{1,0}=\begin{bmatrix}0\\ 1\end{bmatrix}_{1,-\frac{3}{4}}\begin{bmatrix}0\\ 1\end{bmatrix}_{1,-\frac{1}{4}}\begin{bmatrix}0\\ 1\end{bmatrix}_{1,-\frac{1}{4}}\begin{bmatrix}0\\ 1\end{bmatrix}_{1,\frac{3}{4}}\begin{bmatrix}1\\ 0\end{bmatrix}_{1,0}\\ &=\begin{bmatrix}1\\ 0\end{bmatrix}_{1,0}\begin{bmatrix}1\\ 1\end{bmatrix}_{1,-\frac{3}{4}}\begin{bmatrix}1\\ 1\end{bmatrix}_{1,-\frac{1}{4}}\begin{bmatrix}1\\ 1\end{bmatrix}_{1,\frac{1}{4}}\begin{bmatrix}2\\ 3\end{bmatrix}_{1,-\frac{3}{4}}\begin{bmatrix}1\\ 2\end{bmatrix}_{1,-1}\begin{bmatrix}1\\ 2\end{bmatrix}_{1,-\frac{1}{2}}\begin{bmatrix}1\\ 2\end{bmatrix}_{1,0}\begin{bmatrix}1\\ 3\end{bmatrix}_{1,-\frac{3}{4}}\begin{bmatrix}1\\ 1\end{bmatrix}_{1,\frac{3}{4}}\\ &\qquad\times\begin{bmatrix}1\\ 2\end{bmatrix}_{1,0}\begin{bmatrix}1\\ 2\end{bmatrix}_{1,\frac{1}{2}}\begin{bmatrix}1\\ 2\end{bmatrix}_{1,1}\begin{bmatrix}2\\ 5\end{bmatrix}_{1,\frac{3}{4}}\begin{bmatrix}1\\ 3\end{bmatrix}_{1,-\frac{1}{4}}\begin{bmatrix}1\\ 3\end{bmatrix}_{1,\frac{1}{4}}\begin{bmatrix}1\\ 3\end{bmatrix}_{1,\frac{3}{4}}\begin{bmatrix}1\\ 4\end{bmatrix}_{1,0}\begin{bmatrix}0\\ 1\end{bmatrix}_{\frac{1}{4},0}.\end{split}$
Again, this is not yet ordered, and to order it, we need to interchange $[1,3]_{1,-3/4}$ and $[1,1]_{1,3/4}$ in the middle. As the lowest approximation, we consider modulo $G^{>5}$. Then, $\Psi_{1,-3/4}[1,3]$ commutes with other factors, and we have, modulo $G^{>5}$,
(3.83) $\begin{split}&\quad\,\begin{bmatrix}0\\ 1\end{bmatrix}_{\frac{1}{4},0}\begin{bmatrix}1\\ 0\end{bmatrix}_{1,0}\equiv\begin{bmatrix}1\\ 0\end{bmatrix}_{1,0}\begin{bmatrix}1\\ 1\end{bmatrix}_{\frac{1}{4},0}\begin{bmatrix}2\\ 3\end{bmatrix}_{\frac{1}{4},0}\begin{bmatrix}1\\ 2\end{bmatrix}_{\frac{1}{2},-\frac{1}{2}}\begin{bmatrix}1\\ 2\end{bmatrix}_{\frac{1}{2},0}\begin{bmatrix}1\\ 2\end{bmatrix}_{\frac{1}{2},\frac{1}{2}}\begin{bmatrix}1\\ 3\end{bmatrix}_{\frac{1}{4},0}\begin{bmatrix}1\\ 4\end{bmatrix}_{1,0}\begin{bmatrix}0\\ 1\end{bmatrix}_{\frac{1}{4},0}.\end{split}$
To proceed to higher degree, we do in the same way as before. By applying the fission and the pentagon relation (2.25) with $c=2$, we have
(3.84) $\begin{split}&\quad\,\begin{bmatrix}1\\ 3\end{bmatrix}_{1,-\frac{3}{4}}\begin{bmatrix}1\\ 1\end{bmatrix}_{1,\frac{3}{4}}=\begin{bmatrix}1\\ 3\end{bmatrix}_{2,-\frac{7}{4}}\begin{bmatrix}1\\ 3\end{bmatrix}_{2,\frac{1}{4}}\begin{bmatrix}1\\ 1\end{bmatrix}_{2,-\frac{1}{4}}\begin{bmatrix}1\\ 1\end{bmatrix}_{2,\frac{7}{4}}\\ &=\begin{bmatrix}1\\ 1\end{bmatrix}_{2,-\frac{1}{4}}\begin{bmatrix}2\\ 4\end{bmatrix}_{2,-2}\begin{bmatrix}2\\ 4\end{bmatrix}_{2,0}\begin{bmatrix}3\\ 7\end{bmatrix}_{2,-\frac{7}{4}}\begin{bmatrix}1\\ 1\end{bmatrix}_{2,\frac{7}{4}}\begin{bmatrix}2\\ 4\end{bmatrix}_{2,0}\begin{bmatrix}2\\ 4\end{bmatrix}_{2,2}\begin{bmatrix}3\\ 7\end{bmatrix}_{2,\frac{1}{4}}\begin{bmatrix}1\\ 3\end{bmatrix}_{2,-\frac{7}{4}}\begin{bmatrix}1\\ 3\end{bmatrix}_{2,\frac{1}{4}}.\end{split}$
This parallel to (3.38), but with different quantum data. As the next approximation, we consider modulo $G^{>11}$. Then, $[3,7]_{2,-7/4}$ commutes with other factors, and we have, modulo $G^{>11}$,
(3.101) $\begin{bmatrix}1\\ 3\end{bmatrix}_{1,-\frac{3}{4}}\begin{bmatrix}1\\ 1\end{bmatrix}{}_{1,\frac{3}{4}}$ $\equiv\begin{bmatrix}1\\ 1\end{bmatrix}_{1,\frac{3}{4}}\begin{bmatrix}3\\ 5\end{bmatrix}_{1,\frac{3}{4}}\begin{bmatrix}2\\ 4\end{bmatrix}_{1,-1}\begin{bmatrix}2\\ 4\end{bmatrix}_{1,1}\begin{bmatrix}3\\ 7\end{bmatrix}_{1,-\frac{3}{4}}\begin{bmatrix}1\\ 3\end{bmatrix}_{1,-\frac{3}{4}}.$
Then, we plug it into (3.82), and apply the pentagon relation, and we have, modulo $G^{>11}$,
(3.102) $\begin{split}\begin{bmatrix}0\\ 1\end{bmatrix}_{\frac{1}{4},0}\begin{bmatrix}1\\ 0\end{bmatrix}_{1,0}&\equiv\begin{bmatrix}1\\ 0\end{bmatrix}_{1,0}\begin{bmatrix}1\\ 1\end{bmatrix}_{\frac{1}{4},0}\begin{bmatrix}3\\ 4\end{bmatrix}_{1,0}\begin{bmatrix}2\\ 3\end{bmatrix}_{\frac{1}{4},0}\begin{bmatrix}3\\ 5\end{bmatrix}_{\frac{1}{4},0}\begin{bmatrix}4\\ 7\end{bmatrix}_{\frac{1}{4},0}\begin{bmatrix}1\\ 2\end{bmatrix}_{\frac{1}{2},-\frac{1}{2}}\begin{bmatrix}1\\ 2\end{bmatrix}_{\frac{1}{2},0}\begin{bmatrix}1\\ 2\end{bmatrix}_{\frac{1}{2},\frac{1}{2}}\\ &\qquad\times\begin{bmatrix}2\\ 4\end{bmatrix}_{1,-1}\begin{bmatrix}2\\ 4\end{bmatrix}_{1,1}\begin{bmatrix}3\\ 7\end{bmatrix}_{\frac{1}{4},0}\begin{bmatrix}2\\ 5\end{bmatrix}_{\frac{1}{4},0}\begin{bmatrix}3\\ 8\end{bmatrix}_{1,0}\begin{bmatrix}1\\ 3\end{bmatrix}_{\frac{1}{4},0}\begin{bmatrix}1\\ 4\end{bmatrix}_{1,0}\begin{bmatrix}0\\ 1\end{bmatrix}_{\frac{1}{4},0}.\end{split}$
By continuing the procedure modulo $G^{>2^{\ell}-1}$, one can naturally guess that the relation converges to the following one:
(3.103) $\begin{split}\begin{bmatrix}0\\ 1\end{bmatrix}_{\frac{1}{4},0}\begin{bmatrix}1\\ 0\end{bmatrix}_{1,0}&=\begin{bmatrix}1\\ 0\end{bmatrix}_{1,0}\begin{bmatrix}1\\ 1\end{bmatrix}_{\frac{1}{4},0}\begin{bmatrix}3\\ 4\end{bmatrix}_{1,0}\begin{bmatrix}2\\ 3\end{bmatrix}_{\frac{1}{4},0}\begin{bmatrix}5\\ 8\end{bmatrix}_{1,0}\begin{bmatrix}3\\ 5\end{bmatrix}_{\frac{1}{4},0}\cdots\\ &\qquad\times\Biggl{(}\begin{bmatrix}1\\ 2\end{bmatrix}_{\frac{1}{2},0}\prod_{j=0}^{\infty}\begin{bmatrix}2^{j}\\ 2^{j+1}\end{bmatrix}_{2^{j-1},-2^{j-1}}\begin{bmatrix}2^{j}\\ 2^{j+1}\end{bmatrix}_{2^{j-1},2^{j-1}}\Biggr{)}\\ &\qquad\times\cdots\begin{bmatrix}5\\ 12\end{bmatrix}_{1,0}\begin{bmatrix}2\\ 5\end{bmatrix}_{\frac{1}{4},0}\begin{bmatrix}3\\ 8\end{bmatrix}_{1,0}\begin{bmatrix}1\\ 3\end{bmatrix}_{\frac{1}{4},0}\begin{bmatrix}1\\ 4\end{bmatrix}_{1,0}\begin{bmatrix}0\\ 1\end{bmatrix}_{\frac{1}{4},0}.\end{split}$
Again, one can prove the relation (3.103) completely based on the pentagon relation by modifying the proof of [Mat21] for the classical case. The detail will be found in Appendix A.
Note that, in the RHSs of (3.75) and (3.103), all factors have the form $\Psi_{1/\delta(n),b}[n]$.
4. Principal $x$-representation
We introduce another representation of $G$ which we call the principal $x$-representation. It is closely related to the quantization of cluster variables ($x$-variables) in [BZ05]. Then, we obtain the reduction property of wall elements for a QCSD, which is important in our application of the pentagon relation to the positivity problem.
4.1. Principal $x$-representation
Following [GHKK18], we introduce
(4.1) $\tilde{N}=N\oplus M^{\circ},\quad\tilde{M}^{\circ}=M^{\circ}\oplus N,$
which are “dual” to each other twisted by $\delta_{1},\dots,\delta_{r}$. An element of $\tilde{N}$ is denoted by $\tilde{n}=(n,m)$ ($n\in N$, $m\in M^{\circ}$). Similarly, an element of $\tilde{M}^{\circ}$ is denoted by $\tilde{m}=(m,n)$ ($m\in M^{\circ}$, $n\in N$). Below we use the notation such as $\tilde{n}_{1}=(n_{1},m_{1})$, $\tilde{m}^{\prime}=(m^{\prime},n^{\prime})$ without explanation. Meanwhile, we also use the internal-sum notation. For example, for $n\in N$, we write $n=(n,0)\in\tilde{N}$ and $n=(0,n)\in\tilde{M}^{\circ}$. Both spaces $\tilde{N}$ and $\tilde{M}^{\circ}$ are equipped with (essentially common) bilinear forms
(4.2) $\{\tilde{n},\tilde{n}^{\prime}\}$ :=$:=\{n,n^{\prime}\}+\langle n^{\prime},m\rangle-\langle n,m^{\prime}\rangle,$
(4.3) $\{\tilde{m},\tilde{m}^{\prime}\}$ :=$:=\{n,n^{\prime}\}+\langle n^{\prime},m\rangle-\langle n,m^{\prime}\rangle,$
[…]
(4.4) $\langle(n,m),(m^{\prime},n^{\prime})\rangle:=\langle n,m^{\prime}\rangle+\langle n^{\prime},m\rangle.$
Let $\tilde{p}^{*}$ be the group homomorphism defined by
(4.7) $\begin{matrix}\tilde{p}^{*}\colon&N&\rightarrow&\tilde{M}^{\circ}\\ &n&\mapsto&\{\cdot,n\}\end{matrix},\quad\{(n^{\prime},m^{\prime}),n\}=\{n^{\prime},n\}+\langle n,m^{\prime}\rangle.$
Note that the map $\tilde{p}^{*}$ is injective. Let $p^{*}\colon N\rightarrow M^{\circ}$ be the one in Section 3.1. Then, we have
(4.8) $\tilde{p}^{*}(n)=(p^{*}(n),n)\in\tilde{M}^{\circ}.$
The following duality relation holds:
(4.9) $-\{\tilde{p}^{*}(n),\tilde{m}\}=\{\tilde{m},\tilde{p}^{*}(n)\}=\langle n,\tilde{m}\rangle.$
Indeed, by (4.8),
(4.10) $\begin{split}\{\tilde{m},\tilde{p}^{*}(n)\}&=\{(m^{\prime},n^{\prime}),(p^{*}(n),n)\}\\ &=\{n^{\prime},n\}+\langle n,m^{\prime}\rangle-\langle n^{\prime},p^{*}(n)\rangle\\ &=\langle n,m^{\prime}\rangle=\langle n,\tilde{m}\rangle.\end{split}$
Following [GHKK18], let $\tilde{P}\subset\tilde{M}^{\circ}$ be a (not unique) monoid satisfying the following conditions:
  • (i).
    $\tilde{P}=\sigma\cap\tilde{M}^{\circ}$, where $\sigma$ is a $2r$-dimensional strongly convex cone in $\tilde{M}^{\circ}_{\mathbb{R}}$.
  • (ii).
    $\tilde{p}^{*}(e_{1})$, …, $\tilde{p}^{*}(e_{r})\in\tilde{P}$.
Now we are ready to define the (quantum) principal $x$-representation of $G$, which is defined in a parallel way to the $y$-representation.
Let $x$ be a formal symbol. Let $\mathbb{Q}(q^{1/\delta_{0}})[x]_{q}$ be the noncommutative and associative algebra over $\mathbb{Q}(q^{1/\delta_{0}})$ with generators $x^{\tilde{m}}$ ($\tilde{m}\in\tilde{P})$ and the relations
(4.11) $x^{\tilde{m}}x^{\tilde{m}^{\prime}}=q^{-\{\tilde{m},\tilde{m}^{\prime}\}}x^{\tilde{m}+\tilde{m}^{\prime}}.$
Let $\mathcal{R}_{q}(x)$ be the completion of $\mathbb{Q}(q^{1/\delta_{0}})[x]_{q}$ with respect to the maximal ideal generated by $x^{\tilde{m}}$’s ($\tilde{m}\in\tilde{P}\setminus\{0\})$. Any element of $\mathcal{R}_{q}(x)$ is expressed as an infinite sum
(4.12) $\sum_{\tilde{m}\in\tilde{P}}c_{\tilde{m}}x^{\tilde{m}}\quad(c_{\tilde{m}}\in\mathbb{Q}(q^{1/\delta_{0}})).$
We define the action of $\widehat{\mathfrak{g}}$ on $\mathcal{R}_{q}(x)$ by
(4.13) $X_{n}(x^{\tilde{m}}):=$ $\ [\langle n,\tilde{m}\rangle]_{q}x^{\tilde{m}+\tilde{p}^{*}(n)}$
(4.14) $=$ $\ \frac{q^{2\langle n,\tilde{m}\rangle}-1}{q-q^{-1}}x^{\tilde{m}}x^{\tilde{p}^{*}(n)},$
where in the second equality we used the duality relation (4.9). It is easy to check that this is an action of $\widehat{\mathfrak{g}}$ and also a derivation. Thus, it induces the action of $G$ on $\mathcal{R}_{q}(x)$ defined by
(4.15) $(\exp X)(x^{\tilde{m}})=\sum_{j=0}^{\infty}\frac{1}{j!}X^{j}(x^{\tilde{m}})\quad(X\in\widehat{\mathfrak{g}}).$
[…]
Remark 4.1.
[…]
(4.18) $-\Lambda=\begin{pmatrix}O&D\\ -D&DB\end{pmatrix},\quad D=\mathrm{diag}(\delta_{1}^{-1},\cdots,\delta_{r}^{-1}).$
Let $\tilde{B}$ be the principal extension of the matrix $B$,
(4.21) $\tilde{B}=\begin{pmatrix}B\\ I\end{pmatrix}.$
They satisfy the relation
(4.24) $-\Lambda\tilde{B}=\begin{pmatrix}D\\ O\end{pmatrix},$
which is equivalent to the duality relation (4.9). Such a pair $(\Lambda,\tilde{B})$ is called a compatible pair in [BZ05]. The matrix $\Lambda$ was used to define the $q$-commutative relation for $x^{\tilde{m}}$ in [BZ05]. The minus sign for $-\Lambda$ in (4.18) agrees with the minus sign for $-\{\tilde{m},\tilde{m}^{\prime}\}$ in the relation (4.11).
4.2. Action of dilogarithm elements
Let us present an analogous result in Section 2.3 for the principal $x$-representation.
Any $\mathbb{Q}(q^{1/\delta_{0}})$-power of a dilogarithm element $\Psi_{a,b}[n]^{c}$ ($c\in\mathbb{Q}(q^{1/\delta_{0}}))$ acts on $\mathcal{R}_{q}(x)$ under $\rho_{x}^{\mathrm{pr}}$ as
(4.25) $\begin{split}\Psi_{a,b}[n]^{c}(x^{\tilde{m}})&=\exp\Biggl{(}\,c\sum_{j=1}^{\infty}\frac{(-1)^{j+1}}{j[ja]_{q}}q^{jb}X_{jn}\Biggr{)}(x^{\tilde{m}})\\ &=x^{\tilde{m}}\exp\Biggl{(}c\sum_{j=1}^{\infty}\frac{q^{2j\langle n,\tilde{m}\rangle}-1}{q^{2ja}-1}\frac{(-1)^{j+1}}{j}q^{ja}q^{jb}x^{j\tilde{p}^{*}(n)}\Biggr{)}.\end{split}$
Let us consider a parallel situation to (2.20) and (2.21). Fix $n$ in (4.25), and let $a=1/s\delta(n)$, where $s\in\mathbb{Z}_{>0}$ is a divisor of $\delta_{0}/\delta(n)$. Then, $\alpha:=\langle s\delta(n)n,\tilde{m}\rangle$ is an integer, and we have
(4.26) $\begin{split}\frac{q^{2j\langle n,\tilde{m}\rangle}-1}{q^{2ja}-1}&=\frac{q^{2\alpha j/s\delta(n)}-1}{q^{2j/s\delta(n)}-1}=\begin{cases}\displaystyle\sum_{p=0}^{\alpha-1}q^{2jp/s\delta(n)}&\alpha>0,\\ 0&\alpha=0,\\ \displaystyle-\sum_{p=1}^{-\alpha}q^{-2jp/s\delta(n)}&\alpha<0.\end{cases}\end{split}$
Thus, we obtain
(4.27) $\begin{split}&\quad\ \Psi_{1/s\delta(n),b}[n](x^{\tilde{m}})=\begin{cases}\displaystyle x^{\tilde{m}}\prod_{p=1}^{\alpha}(1+q^{(2p-1)/s\delta(n)}q^{b}x^{\tilde{p}^{*}(n)})&\alpha>0,\\ x^{\tilde{m}}&\alpha=0,\\ \displaystyle x^{\tilde{m}}\prod_{p=1}^{-\alpha}(1+q^{-(2p-1)/s\delta(n)}q^{b}x^{\tilde{p}^{*}(n)})^{-1}&\alpha<0.\end{cases}\end{split}$
Observe that this is indeed the automorphism part of the Fock-Goncharov decomposition of mutations of the quantum $x$-variables with principal coefficients in [BZ05]. See also [Man21, DM21, CFMM20] for an alternative approach, where the same representation is described through the adjoint action of the quantum dilogarithm on $\mathcal{R}_{q}(x)$ as in Remark 2.6.
4.3. Reduction of wall elements in a QCSD
Let $\mathcal{S}_{q}(x)$ be the subset of $\mathcal{R}_{q}(x)$ where the coefficients $c_{\tilde{m}}$ in (4.12) are restricted in $\mathbb{Z}[q^{\pm 1/\delta_{0}}]$. Then, $\mathcal{S}_{q}(x)$ is an algebra over $\mathbb{Z}[q^{\pm 1/\delta_{0}}]$ by the same relation (4.11). The following lemma is a quantum analogue of the one in [GHKK18, Appendix C.3 Step II], [Nak23, Lemma III.5.8].
Lemma 4.2.
A $\mathbb{Q}(q^{1/\delta_{0}})$-power of a quantum dilogarithm element $\Psi_{a,b}[n]^{c}$ $(c\in\mathbb{Q}(q^{1/\delta_{0}}))$ acts on $\mathcal{S}_{q}(x)$ if and only if $a=1/s\delta(n)$ for a divisor $s$ of $\delta_{0}/\delta(n)$ and $c\in\mathbb{Z}$.
Proof.
The if-part is true by (4.27). Let us prove the only-if-part. Assume that $\Psi_{a,b}[n]^{c}$ acts on $\mathcal{S}_{q}(x)$. Then, in the second expression of the RHS of (4.25), the term
(4.28) $c\frac{q^{2j\langle n,\tilde{m}\rangle}-1}{q^{2ja}-1}\frac{(-1)^{j+1}}{j}q^{ja}q^{jb}$
should at least belong to $\mathbb{Q}[q^{\pm 1/\delta_{0}}]$ for any $j$. Let $a=1/s\delta(n)$, where $s$ is a positive rational number. Then, the condition implies $s\langle\delta(n)n,\tilde{m}\rangle\in\mathbb{Z}$ for any $\tilde{m}\in\tilde{P}$. By the assumption for $\tilde{P}$, $\tilde{P}$ contains a basis of $\tilde{M}^{\circ}$. It follows that $s\langle\delta(n)n,\tilde{m}\rangle\in\mathbb{Z}$ for any $\tilde{m}\in\tilde{M}^{\circ}$. Due to the definition of $\delta(n)$, there is some $\tilde{m}\in\tilde{M}^{\circ}$ such that $\langle\delta(n)n,\tilde{m}\rangle=1$. This implies $s\in\mathbb{Z}$. On the other hand, let $t=\delta_{0}/\delta(n)\in\mathbb{Z}_{>0}$. Then, $a=t/s\delta_{0}$. Since we require that $a\in(1/\delta_{0})\mathbb{Z}_{>0}$, $s$ is a divisor of $t$. Finally, by (4.27) and the assumption, we have $c\in\mathbb{Z}$. ∎
We have the following important corollary of Lemma 4.2.
Proposition 4.3.
Any QCSD $\mathfrak{D}_{\mathfrak{s}}^{q}$ is equivalent to a consistent scattering diagram whose wall elements have the form
(4.29) $\Psi_{1/\delta(n),b}[n]^{\pm c}\quad(n\in N^{+},\ b\in(1/\delta_{0})\mathbb{Z},\ c\in\mathbb{Z}).$
Proof.
This can be proved inductively on the degree of $n$ as in the classical case by considering the consistency relations around the perpendicular joints. See Step 2 in the proof of [GHKK18, Prop. C.13]. For the reader’s convenience, we give the key point of the inductive step. Suppose that a scattering diagram have the walls whose wall elements have the form (4.29) with $\deg n\leq\ell$ and that, for a small loop around a perpendicular joint $\gamma$, the consistency relation around $\gamma$ holds modulo $G^{>\ell}$. By the if-part of Lemma 4.2, all wall elements act on $\mathcal{S}_{q}(x)$. Then, we may add walls whose wall elements have the form (4.29) so that the consistency relation around $\gamma$ holds modulo $G^{>\ell+1}$. ∎
Example 4.4.
All examples of QCSDs in Section 3 have the wall elements of the form (4.29). Moreover, there are no negative powers $\Psi_{1/\delta(n),b}[n]^{-1}$.
5. Application to positivity and nonpositivity of QCSDs
In the rest of the paper we apply the pentagon relation, together with the combinatorics of quantum data, to study the positivity and nonpositivity property of QCSDs.
5.1. Positivity of QCSDs
Let us briefly summarize known results concerning the positivity of CSDs and QCSDs. Below, “skew-symmetric” means the initial exchange matrix in (2.1) is skew-symmetric.
  1. (1)
    Positivity of theta functions of CSD [GHKK18]. For any CSD (including nonskew-symmetric one), nonzero coefficients of any theta function are positive integers (positivity). In addition, nonzero multiplicities of the product of elements of theta basis are positive integers (strong positivity).
  2. (2)
    Positivity of theta functions of QCSD [DM21]. For any skew-symmetric QCSD, nonzero coefficients of any (quantum) theta function belong to $\mathbb{Z}_{\geq 0}[q^{\pm 1/\delta_{0}}]$ (positivity). In addition, nonzero multiplicities of the product of elements of theta basis belong to $\mathbb{Z}_{\geq 0}[q^{\pm 1/\delta_{0}}]$ (strong positivity).
  3. (3)
    Positivity and nonpositivity of quantum greedy basis [LLRZ14]. Consider a cluster algebra of rank 2 with the initial exchange matrix
    (5.3) $B=\begin{pmatrix}0&-\delta_{1}\\ \delta_{2}&0\end{pmatrix}\quad(\delta_{1},\delta_{2}\in\mathbb{Z}_{>0}).$
    Then, the positivity of the quantum greedy elements fails for $(\delta_{1},\delta_{2})\allowbreak=(2,3)$, $(2,5)$, $(3,4)$, $(4,6)$, for example. Also, it was conjectured [LLRZ14, Conj. 14] that the strong positivity holds for the quantum greedy basis if $\delta_{1}|\delta_{2}$ or $\delta_{2}|\delta_{1}$. It is known that in the classical case the greedy basis and the theta basis coincide [${}^{+}$17]. Also, the above nonpositivity of the quantum greedy elements was explained in view of the nonpositivity of the wall functions in the corresponding QCSD [CFMM20] for the case $(\delta_{1},\delta_{2})=(2,3)$.
We see in the above that there is an intriguing discrepancy of the positivity property between CSDs and QCSDs especially in the nonskew-symmetric case. Then, we have the following natural questions.
  1. Q1:
    Exactly when and why does the positivity property of a nonskew-symmetric QCSD hold or fail?
  2. Q2:
    When the positivity property fails for a QCSD, how is the positivity property restored in the classical limit $q\rightarrow 1$?
Based on various results by [GHKK18, Nak23, Man21, DM21, LLRZ14, ${}^{+}$17, CFMM20], and also in view of Proposition 4.3, we formulate the positivity of QCSD as follows:
Definition 5.1.
A QCSD $\mathfrak{D}_{\mathfrak{s}}^{q}$ is a positive realization if the wall element of any wall has the form
(5.4) $\Psi_{1/\delta(n),b}[n]^{c}\quad(n\in N^{+},\ b\in(1/\delta_{0})\mathbb{Z},\ c\in\mathbb{Z}_{>0}).$
Namely, the negative powers $\Psi_{1/\delta(n),b}[n]^{-c}$ ($c\in\mathbb{Z}_{>0}$) do not appear as wall elements. We say that a QCSD $\mathfrak{D}_{\mathfrak{s}}^{q}$ is positive if it is equivalent to a consistent scattering diagram that is a positive realization.
Example 5.2.
All examples of QCSDs in Section 3 are positive realizations, thus, positive as observed in Example 4.4.
For a QCSD, a theta function is defined by broken lines [DM21] as in the classical one [GHKK18].
Proposition 5.3 ([DM21, Theorem 3.9]).
Every theta function of a QCSD $\mathfrak{D}_{\mathfrak{s}}^{q}$ has a formal power series of the form
(5.5) $\theta_{Q,\tilde{m}_{0}}=x^{\tilde{m}_{0}}\biggl{(}1+\sum_{n\in N^{+}}c_{n}x^{\tilde{p}^{*}(n)}\biggr{)}\quad(c_{n}\in\mathbb{Z}_{\geq 0}[q^{\pm 1/\delta_{0}}]).$
if $\mathfrak{D}_{\mathfrak{s}}^{q}$ is positive.
This was proved in [DM21, Theorem 3.9] by the same method for the classical case [GHKK18, Theorem 1.13, Remark 3.2]. Similarly, the positivity of a QCSD $\mathfrak{D}_{\mathfrak{s}}^{q}$ also implies the strong positivity of the theta basis [DM21, Theorem 3.10] as in the classical case [GHKK18].
We also expect that the converse of Proposition 5.3 holds. Let us concentrate on a particular situation, which is relevant to the forthcoming Theorem 5.6.
Example 5.4.
Consider the rank 2 case. Assume that there is some outgoing wall $\mathbf{w}$ in $\mathfrak{D}_{\mathfrak{s}}^{q}$ whose normal vector is $n=(1,1)$ and whose wall element has the form $\Psi_{1/\delta(n),b}[n]^{-c}$ ($c\in\mathbb{Z}_{>0}$), and it is not cancelled by the same type of walls with positive power. Take a point $Q$ in the same quadrant with $\mathbf{w}$. Consider a theta function $\theta_{Q,\tilde{m}_{0}}$ to which a broken line $\gamma$ bending only at $\mathbf{w}$ contributes. Then, $\gamma$ may carry a coefficient $c_{\gamma}$ not in $\mathbb{Z}_{\geq 0}[q^{\pm 1/\delta_{0}}]$. Also, the contribution of $\gamma$ to $\theta_{Q,\tilde{m}_{0}}$ is not cancelled from any other broken lines. Indeed, if there is such a broken line $\gamma^{\prime}$, it should bend on both of the two incoming walls and ends at $Q$. Thus, $\gamma^{\prime}$ should encircle the origin completely. But, this is impossible due to the definition of broken lines. Therefore, the converse of Proposition 5.3 holds in this case.
As already mentioned, the following result is known.
Theorem 5.5 ([DM21, Theorem 2.15]).
Every skew-symmetric QCSD $\mathfrak{D}_{\mathfrak{s}}^{q}$ is positive.
The proof of [DM21] is based on the Donaldson-Thomas theory on quiver-representations, which is applicable only to the skew-symmetric case. Below we study the positivity of the nonskew-symmetric case by means of the pentagon relation.
5.2. Rank 2 examples
Let us consider the rank 2 case. As mentioned in Example 5.2, all examples with $\delta_{1}\delta_{2}\leq 4$ in Section 3 are positive. Let us consider examples with $\delta_{1}\delta_{2}>4$, namely, QCSDs of nonaffine infinite type. We use the same conventions and notations in Section 3.
Let us compute in the lowest approximation, namely, modulo $G^{>2}$.
(a). $(\delta_{1},\delta_{2})=(2,3)$. This is the case where the nonpositivity of a quantum greedy element is presented in [LLRZ14], and further examined in view of the QCSD in [CFMM20]. We have
(5.6) $\begin{split}\begin{bmatrix}0\\ 1\end{bmatrix}_{\frac{1}{3},0}\begin{bmatrix}1\\ 0\end{bmatrix}_{\frac{1}{2},0}&\equiv\begin{bmatrix}1\\ 0\end{bmatrix}_{\frac{1}{2},0}\begin{bmatrix}1\\ 1\end{bmatrix}_{1,-\frac{7}{6}}\begin{bmatrix}1\\ 1\end{bmatrix}_{1,-\frac{1}{2}}\begin{bmatrix}1\\ 1\end{bmatrix}_{1,-\frac{1}{6}}\begin{bmatrix}1\\ 1\end{bmatrix}_{1,\frac{1}{6}}\begin{bmatrix}1\\ 1\end{bmatrix}_{1,\frac{1}{2}}\begin{bmatrix}1\\ 1\end{bmatrix}_{1,\frac{7}{6}}\begin{bmatrix}0\\ 1\end{bmatrix}_{\frac{1}{3},0}\\ &=\begin{bmatrix}1\\ 0\end{bmatrix}_{\frac{1}{2},0}\begin{bmatrix}1\\ 1\end{bmatrix}_{\frac{1}{6},-\frac{1}{3}}\begin{bmatrix}1\\ 1\end{bmatrix}_{\frac{1}{6},0}^{-1}\begin{bmatrix}1\\ 1\end{bmatrix}_{\frac{1}{6},\frac{1}{3}}\begin{bmatrix}0\\ 1\end{bmatrix}_{\frac{1}{3},0}.\end{split}$
Thus, we confirmed again that it is nonpositive in the sense of Definition 5.1. Moreover, in the classical limit, the negative power in the last expression is cancelled by the positive power, and we obtain $\Psi[\mathbf{e}_{1}]^{2}\Psi[(1,1)]^{6}\Psi[\mathbf{e}_{2}]^{3}$. This clarifies the mechanism how the positivity is restored in the classical limit. Thus, we have an answer to Question 2 in Section 5.1. Moreover, comparing the computation in (5.6) with the much more complicated one in [CFMM20, Appendix B], we observe that the pentagon relation significantly simplifies the analysis (just in two lines in (5.6)).
(b). $(\delta_{1},\delta_{2})=(2,4)$. By a similar computation, we have
(5.7) $\begin{split}\begin{bmatrix}0\\ 1\end{bmatrix}_{\frac{1}{4},0}\begin{bmatrix}1\\ 0\end{bmatrix}_{\frac{1}{2},0}\equiv\begin{bmatrix}1\\ 0\end{bmatrix}_{\frac{1}{2},0}\begin{bmatrix}1\\ 1\end{bmatrix}_{\frac{1}{4},-\frac{1}{2}}\begin{bmatrix}1\\ 1\end{bmatrix}_{\frac{1}{4},\frac{1}{2}}\begin{bmatrix}0\\ 1\end{bmatrix}_{\frac{1}{4},0}.\end{split}$
This is positive up to $G^{\leq 2}$.
(c). $(\delta_{1},\delta_{2})=(3,3)$. This is skew-symmetric, thus, positive by Theorem 5.5. Indeed, we have
(5.8) $\begin{split}\begin{bmatrix}0\\ 1\end{bmatrix}_{\frac{1}{3},0}\begin{bmatrix}1\\ 0\end{bmatrix}_{\frac{1}{3},0}\equiv\begin{bmatrix}1\\ 0\end{bmatrix}_{\frac{1}{3},0}\begin{bmatrix}1\\ 1\end{bmatrix}_{\frac{1}{3},-\frac{2}{3}}\begin{bmatrix}1\\ 1\end{bmatrix}_{\frac{1}{3},0}\begin{bmatrix}1\\ 1\end{bmatrix}_{\frac{1}{3},\frac{2}{3}}\begin{bmatrix}0\\ 1\end{bmatrix}_{\frac{1}{3},0}.\end{split}$
This is positive up to $G^{\leq 2}$ as it should be.
(d). $(\delta_{1},\delta_{2})=(3,4)$. This is similar to Example (a). We have
(5.9) $\begin{split}\begin{bmatrix}0\\ 1\end{bmatrix}_{\frac{1}{3},0}\begin{bmatrix}1\\ 0\end{bmatrix}_{\frac{1}{4},0}\equiv\begin{bmatrix}1\\ 0\end{bmatrix}_{\frac{1}{4},0}\begin{bmatrix}1\\ 1\end{bmatrix}_{\frac{1}{12},-\frac{1}{2}}\begin{bmatrix}1\\ 1\end{bmatrix}_{\frac{1}{12},-\frac{1}{3}}^{-1}\begin{bmatrix}1\\ 1\end{bmatrix}_{\frac{1}{12},0}\begin{bmatrix}1\\ 1\end{bmatrix}_{\frac{1}{12},\frac{1}{3}}^{-1}\begin{bmatrix}1\\ 1\end{bmatrix}_{\frac{1}{12},\frac{1}{2}}\begin{bmatrix}0\\ 1\end{bmatrix}_{\frac{1}{3},0}.\end{split}$
This is nonpositive.
5.3. Nonpositivity result
Consider a cluster algebra of rank 2 such that its exchange matrix $B$ has the form in (5.3). In [LLRZ14, Conj. 14], it was conjecture that the strong positivity of the quantum greedy basis holds if $\delta_{1}|\delta_{2}$ or $\delta_{2}|\delta_{1}$. The following result is motivated by this conjecture.
Theorem 5.6.
Let $\mathfrak{D}^{q}_{\mathfrak{s}}$ be a QCSD of rank 2 such that $b_{12}\neq 0$ for $B$ in (2.1). Then, $\mathfrak{D}^{q}_{\mathfrak{s}}$ is nonpositive if $\delta_{1}\nmid\delta_{2}$ and $\delta_{2}\nmid\delta_{1}$.
Proof.
We show that under the condition the nonpositivity already emerges at degree 2 just as above Examples (a) and (d). We continue to assume that $\{e_{2},e_{1}\}=1$. By a similar consideration to the examples in Section 5.2, we have, modulo $G^{>2}$,
(5.10) $\begin{split}\begin{bmatrix}0\\ 1\end{bmatrix}_{\frac{1}{\delta_{2}},0}\begin{bmatrix}1\\ 0\end{bmatrix}_{\frac{1}{\delta_{1}},0}&=\Biggl{(}\prod_{j=1}^{\delta_{2}}\begin{bmatrix}0\\ 1\end{bmatrix}_{1,\frac{\delta_{2}+1-2j}{\delta_{2}}}\Biggr{)}\Biggl{(}\prod_{i=1}^{\delta_{1}}\begin{bmatrix}1\\ 0\end{bmatrix}_{1,\frac{\delta_{1}+1-2i}{\delta_{1}}}\Biggr{)}\\ &\equiv\begin{bmatrix}1\\ 0\end{bmatrix}_{\frac{1}{\delta_{1}},0}\Biggl{(}\prod_{i=1}^{\delta_{1}}\prod_{j=1}^{\delta_{2}}\begin{bmatrix}1\\ 1\end{bmatrix}_{1,b(i,j)}\Biggr{)}\begin{bmatrix}0\\ 1\end{bmatrix}_{\frac{1}{\delta_{2}},0},\end{split}$
(5.11) $b(i,j)=(2\delta_{1}\delta_{2}+(1-2j)\delta_{1}+(1-2i)\delta_{2})/\delta_{1}\delta_{2}.$
Let $c=\mathrm{gcd}(\delta_{1},\delta_{2})$, and let $\delta_{1}=\delta^{\prime}_{1}c$, $\delta_{2}=\delta^{\prime}_{2}c$. Then, $\delta((1,1))=\delta^{\prime}_{1}\delta^{\prime}_{2}c$, and we have
(5.12) $b(i,j)=(2\delta^{\prime}_{1}\delta^{\prime}_{2}c+\delta^{\prime}_{1}+\delta^{\prime}_{2}-2\delta^{\prime}_{1}j-2\delta^{\prime}_{2}i)/\delta((1,1)).$
Let $m_{0}$ be the largest number of the numerator of $b(i,j)$, which is attained with $i=j=1$. The exponents of $q$ in $\Psi_{1/\delta((1,1)),b}[(1,1)]$ align with the interval $2/\delta((1,1))$. Thus, if the positivity holds, we should also have the number $m_{0}-2$ in the numerator. This occurs only when $\delta^{\prime}_{1}=1$ or $\delta^{\prime}_{2}=1$. This means $\delta_{1}\mid\delta_{2}$ or $\delta_{2}\mid\delta_{1}$. ∎
Note that this is exactly the situation in Example 5.4.
We have the following corollary of Theorem 5.6, where we convert the statement into a necessary condition for the positivity.
Corollary 5.7.
Let $\mathfrak{D}^{q}_{\mathfrak{s}}$ be a QCSD of any rank. Then, if $\mathfrak{D}^{q}_{\mathfrak{s}}$ is positive, $\delta_{i}\mid\delta_{j}$ or $\delta_{j}\mid\delta_{i}$ holds for any pair $i\neq j$ such that $b_{ij}\neq 0$ for $B$ in (2.1).
Proof.
Suppose that $\delta_{i}\nmid\delta_{j}$ holds for some $i$ and $j$ such that $b_{ij}\neq 0$. We consider a QCSD whose wall elements are in the form (4.29). Then, the proof of Theorem 5.6 tells that there is some wall whose wall element is $\Psi_{1/\delta(e_{i}+e_{j}),b}[e_{i}+e_{j}]^{-1}$ for some $b$. Moreover, the wall element $\Psi_{1/\delta(e_{i}+e_{j}),b}[e_{i}+e_{j}]$ arises only from the situation in the proof of Theorem 5.6 due to the construction of a CSD in [GHKK18, Appendix C.1]. Therefore, the wall element $\Psi_{1/\delta(e_{i}+e_{j}),b}[e_{i}+e_{j}]^{-1}$ is never cancelled by other wall elements. ∎
5.4. Positivity up to degree 4
We are left with the problem to determine whether the converses of Theorem 5.6 and Corollary 5.7 hold or not. Here, we concentrate on the former one and examine the positivity up to some degrees.
First, let us give some examples.
Example 5.8.
(a). Let $(\delta_{1},\delta_{2})=(1,5)$. By a direct computation with the pentagon relation, we verify the following relation modulo $G^{>6}$:
(5.13) $\begin{split}\begin{bmatrix}0\\ 1\end{bmatrix}_{\frac{1}{5},0}\begin{bmatrix}1\\ 0\end{bmatrix}_{1,0}&\equiv\begin{bmatrix}1\\ 0\end{bmatrix}_{1,0}\begin{bmatrix}1\\ 1\end{bmatrix}_{\frac{1}{5},0}\begin{bmatrix}2\\ 3\end{bmatrix}_{\frac{1}{5},-\frac{2}{5}}\begin{bmatrix}2\\ 3\end{bmatrix}_{\frac{1}{5},\frac{2}{5}}\begin{bmatrix}1\\ 2\end{bmatrix}_{\frac{1}{5},-\frac{2}{5}}\begin{bmatrix}1\\ 2\end{bmatrix}_{\frac{1}{5},\frac{2}{5}}\\ &\qquad\times\begin{bmatrix}2\\ 4\end{bmatrix}_{\frac{2}{5},-\frac{6}{5}}\begin{bmatrix}2\\ 4\end{bmatrix}_{\frac{2}{5},-\frac{4}{5}}\begin{bmatrix}2\\ 4\end{bmatrix}_{\frac{2}{5},\frac{4}{5}}\begin{bmatrix}2\\ 4\end{bmatrix}_{\frac{2}{5},\frac{6}{5}}\\ &\qquad\times\begin{bmatrix}1\\ 3\end{bmatrix}_{\frac{1}{5},-\frac{2}{5}}\begin{bmatrix}1\\ 3\end{bmatrix}_{\frac{1}{5},\frac{2}{5}}\begin{bmatrix}1\\ 4\end{bmatrix}_{\frac{1}{5},0}\begin{bmatrix}1\\ 5\end{bmatrix}_{1,0}\begin{bmatrix}0\\ 1\end{bmatrix}_{\frac{1}{5},0}.\end{split}$
Thus, it is positive up to $G^{\leq 6}$.
(b). Let $(\delta_{1},\delta_{2})=(2,4)$. We have the following relation modulo $G^{>4}$:
(5.14) $\begin{split}\begin{bmatrix}0\\ 1\end{bmatrix}_{\frac{1}{4},0}\begin{bmatrix}1\\ 0\end{bmatrix}_{\frac{1}{2},0}&\equiv\begin{bmatrix}1\\ 0\end{bmatrix}_{\frac{1}{2},0}\begin{bmatrix}2\\ 1\end{bmatrix}_{\frac{1}{4},0}\begin{bmatrix}1\\ 1\end{bmatrix}_{\frac{1}{4},-\frac{1}{2}}\begin{bmatrix}1\\ 1\end{bmatrix}_{\frac{1}{4},\frac{1}{2}}\begin{bmatrix}2\\ 2\end{bmatrix}_{\frac{1}{2},-\frac{3}{2}}\begin{bmatrix}2\\ 2\end{bmatrix}_{\frac{1}{2},-1}\\ &\qquad\times\begin{bmatrix}2\\ 2\end{bmatrix}_{\frac{1}{2},-\frac{1}{2}}\begin{bmatrix}2\\ 2\end{bmatrix}_{\frac{1}{2},\frac{1}{2}}\begin{bmatrix}2\\ 2\end{bmatrix}_{\frac{1}{2},1}\begin{bmatrix}2\\ 2\end{bmatrix}_{\frac{1}{2},\frac{3}{2}}\begin{bmatrix}1\\ 2\end{bmatrix}_{\frac{1}{2},-1}\begin{bmatrix}1\\ 2\end{bmatrix}_{\frac{1}{2},-\frac{1}{2}}\\ &\qquad\times\begin{bmatrix}1\\ 2\end{bmatrix}_{\frac{1}{2},0}\begin{bmatrix}1\\ 2\end{bmatrix}_{\frac{1}{2},0}\begin{bmatrix}1\\ 2\end{bmatrix}_{\frac{1}{2},\frac{1}{2}}\begin{bmatrix}1\\ 2\end{bmatrix}_{\frac{1}{2},1}\begin{bmatrix}1\\ 3\end{bmatrix}_{\frac{1}{4},-\frac{1}{2}}\begin{bmatrix}1\\ 3\end{bmatrix}_{\frac{1}{4},\frac{1}{2}}\begin{bmatrix}0\\ 1\end{bmatrix}_{\frac{1}{4},0}.\end{split}$
Thus, it is positive up to $G^{\leq 4}$.
We can do the same computation up to $G^{\leq 4}$ in full generality and confirm the positivity as below.
Example 5.9.
Suppose that $\delta_{2}=k\delta_{1}$ for some positive integer $k$. Applying the pentagon relation and the fission/fusion formula as before, we obtain the following explicit formulas of the factors in the ordered product for the LHS of (5.10) up to degree 4.
(a). Degree 2. We have $\delta((1,1))=k\delta_{1}$, and the formula in (5.10) is specialized as
(5.19) $\prod_{j=1}^{k\delta_{1}}\prod_{i=1}^{\delta_{1}}\begin{bmatrix}1\\ 1\end{bmatrix}_{1,\frac{2k\delta_{1}+1+k-2j-2ki}{k\delta_{1}}}=\prod_{i=1}^{\delta_{1}}\begin{bmatrix}1\\ 1\end{bmatrix}_{\frac{1}{k\delta_{1}},\frac{\delta_{1}+1-2i}{\delta_{1}}}.$
(b). Degree 3. We have
(5.20) $\delta((2,1))=k\delta_{1},\quad\delta((1,2))=\begin{cases}k\delta_{1}/2&(k:\text{even}),\\ k\delta_{1}&(k:\text{odd}).\end{cases}$
Then, we have the terms for $n=(2,1)$ and $(1,2)$,
(5.25) $\prod_{j=1}^{k\delta_{1}}\prod_{i=1}^{\delta_{1}-1}\prod_{t=1}^{i}\begin{bmatrix}2\\ 1\end{bmatrix}_{1,\frac{-k\delta_{1}+1-2j+2ki+2kt}{k\delta_{1}}}=\prod_{i=1}^{\delta_{1}-1}\prod_{t=1}^{i}\begin{bmatrix}2\\ 1\end{bmatrix}_{\frac{1}{k\delta_{1}},\frac{-2\delta_{1}+2i+2t}{\delta_{1}}},$
(5.26) $\begin{split}\prod_{j=1}^{k\delta_{1}-1}\prod_{i=1}^{\delta_{1}}\prod_{t=1}^{k\delta_{1}-j}\begin{bmatrix}1\\ 2\end{bmatrix}_{1,\frac{k\delta_{1}+k-2j-2ki+2t}{k\delta_{1}}}\hskip 100.0pt\\ =\begin{cases}\displaystyle\prod_{i=1}^{\delta_{1}}\prod_{t=1}^{k\delta_{1}-1}\begin{bmatrix}1\\ 2\end{bmatrix}_{\frac{2}{k\delta_{1}},\frac{k-2ki+2t}{k\delta_{1}}}&(k\delta_{1}:\text{even}),\\ \displaystyle\prod_{i=1}^{\delta_{1}}\prod_{t=1}^{(k\delta_{1}-1)/2}\begin{bmatrix}1\\ 2\end{bmatrix}_{\frac{1}{k\delta_{1}},\frac{k-1-2ki+4t}{k\delta_{1}}}&(k\delta_{1}:\text{odd}).\end{cases}\end{split}$
The equality (5.25) is straightforward, while the one (5.26) follows from the following resummation formula for any positive integer $p$:
(5.27) $\begin{split}\sum_{j=1}^{p-1}\sum_{t=1}^{p-j}x^{-2j+2t}&=\frac{x^{-2p+4}(1-x^{2p-2})(1-x^{2p})}{(1-x^{2})(1-x^{4})}\\ &=\begin{cases}\displaystyle\sum_{j=1}^{p/2}\sum_{t=1}^{p-1}x^{2-4j+2t}&(p:\text{even}),\\ \displaystyle\sum_{j=1}^{p}\sum_{t=1}^{(p-1)/2}x^{-2j+4t}&(p:\text{odd}).\end{cases}\end{split}$
To see the positivity, we need to resolve the discrepancy between (5.20) and (5.26) for $n=(1,2)$ when $\delta_{1}$ is even and $k$ is odd, e.g., $(\delta_{1},\delta_{2})=(2,6)$. In fact, in this case we can further rewrite the RHS of (5.26) (for even $k\delta_{1}$) as
(5.28) $\begin{split}&\prod_{i=1}^{\delta_{1}/2}\Biggl{\{}\Biggl{(}\prod_{t=1}^{(k+1)/2}\begin{bmatrix}1\\ 2\end{bmatrix}_{\frac{1}{k\delta_{1}},\frac{k-1-4ki+4t}{k\delta_{1}}}\begin{bmatrix}1\\ 2\end{bmatrix}_{\frac{1}{k\delta_{1}},\frac{k+2k\delta_{1}-5-4ki+4t}{k\delta_{1}}}\Biggr{)}\prod_{t=1}^{k\delta_{1}-k-2}\begin{bmatrix}1\\ 2\end{bmatrix}_{\frac{1}{k\delta_{1}},\frac{3k+1-4ki+2t}{k\delta_{1}}}\Biggr{\}}.\end{split}$
(c). Degree 4. The results are similar to the degree 3 case. However, the formulas become lengthy, so that we put them in Appendix B. They also imply the difficulty to prove the positivity in higher degrees by direct computation.
We summarize the above results in Example 5.9 as follows.
Proposition 5.10.
The converse of Theorem 5.6 holds at least up to $G^{\leq 4}$.
Encouraged by this partial but already nontrivial result, we give a conjecture in the spirit of [LLRZ14, Conj. 14].
Conjecture 5.11 (Cf. [LLRZ14, Conj. 14]).
The converse of Theorem 5.6 holds.
Appendix A Derivation of relations (3.75) and (3.103) by pentagon relation
We faithfully follow the derivation of [Mat21] in the classical case. We ask the reader to consult [Mat21] for further details. It is important that in the derivation we only use the pentagon and commutative relations in Theorem 2.7.
A.1. Derivation of relation (3.75)
Let us write $\Psi_{a,b}[n]$ as $[n]_{a,b}$ for simplicity.
Lemma A.1 (cf. [Mat21, Lemma 1]).
If $\{n^{\prime},n\}=c>0$, the following formulas hold:
(A.1) $[n^{\prime}]_{c/2,b^{\prime}}[n]_{c,b}$ $=[n]_{c,b}[n+n^{\prime}]_{c/2,b+b^{\prime}}[n+2n^{\prime}]_{c,b+2b^{\prime}}[n^{\prime}]_{c/2,b^{\prime}},$
(A.2) $[n^{\prime}]_{c,b^{\prime}}[n]_{c/2,b}$ $=[n]_{c/2,b}[2n+n^{\prime}]_{c,2b+b^{\prime}}[n+n^{\prime}]_{c/2,b+b^{\prime}}[n^{\prime}]_{c,b^{\prime}}.$
Proof.
This is proved by the fission/fusion and the pentagon relation in the same way as the relation (3.27). ∎
Lemma A.2 (cf. [Mat21, Lemma 2]).
Let $\ell$ be any nonnegative integer. If $\{n^{\prime},n\}=c>0$, the following formula holds:
(A.3) $\begin{split}&\ [n^{\prime}]_{c/2,b^{\prime}}\biggl{(}\,\buildrel\longrightarrow\over{\prod_{0\leq p\leq\ell}}[n+2pn^{\prime}]_{c,b+2pb^{\prime}\pm pc}\biggr{)}\\ =&\ [n]_{c,b}\biggl{(}\,\buildrel\longrightarrow\over{\prod_{1\leq p\leq 2\ell+1}}[n+pn^{\prime}]_{c/2,b+pb^{\prime}\pm(p-1)c/2}\biggr{)}[n+(2\ell+2)n^{\prime}]_{c,b+(2\ell+2)b^{\prime}\pm\ell c}[n^{\prime}]_{c/2,b^{\prime}},\end{split}$
(A.4) $\begin{split}&\ \biggl{(}\,\buildrel\longleftarrow\over{\prod_{0\leq p\leq\ell}}[2pn+n^{\prime}]_{c,2pb+b^{\prime}\pm pc}\biggr{)}[n]_{c/2,b}\\ =&\ [n]_{c/2,b}[(2\ell+2)n+n^{\prime}]_{c,(2\ell+2)b+b^{\prime}\pm\ell c}\biggl{(}\,\buildrel\longleftarrow\over{\prod_{1\leq p\leq 2\ell+1}}[pn+n^{\prime}]_{c/2,pb+b^{\prime}\pm(p-1)c/2}\biggr{)}[n^{\prime}]_{c,b^{\prime}}.\end{split}$
Proof.
For $\ell=0$, they coincide with the relations in Lemma A.1. Then, they are proved by the induction on $\ell$ with the pentagon relation and Lemma A.1 in the same way as [Mat21, Lemma 2]. ∎
Lemma A.3 (cf. [Mat21, Lemma 3]).
Let $\ell$ be any nonnegative integer. If $\{n^{\prime},n\}=c>0$, the following formulas hold:
(A.5) $\begin{split}[n^{\prime}]_{c/2,b^{\prime}}\biggl{(}\,\buildrel\longrightarrow\over{\prod_{p\geq 0}}[n+2pn^{\prime}]_{c,b+2pb^{\prime}\pm pc}\biggr{)}&=[n]_{c,b}\biggl{(}\,\buildrel\longrightarrow\over{\prod_{p\geq 1}}[n+pn^{\prime}]_{c/2,b+pb^{\prime}\pm(p-1)c/2}\biggr{)}[n^{\prime}]_{c/2,b^{\prime}},\end{split}$
(A.6) $\begin{split}\biggl{(}\,\buildrel\longleftarrow\over{\prod_{p\geq 0}}[2pn+n^{\prime}]_{c,2pb+b^{\prime}\pm pc}\biggr{)}[n]_{c/2,b}&=[n]_{c/2,b}\biggl{(}\,\buildrel\longleftarrow\over{\prod_{p\geq 1}}[pn+n^{\prime}]_{c/2,pb+b^{\prime}\pm(p-1)c/2}\biggr{)}[n^{\prime}]_{c,b^{\prime}}.\end{split}$
Proof.
This is obtained by taking the limit $\ell\rightarrow\infty$ of the relations in Lemma A.2. ∎
Theorem A.4 (cf. [Mat21, Theorem 2]).
If $\{n^{\prime},n\}=c>0$, the following formula holds:
(A.7) $\begin{split}&\ [n^{\prime}]_{c/2,b^{\prime}}[n]_{c/2,b}\\ =&\ \biggl{(}\,\buildrel\longrightarrow\over{\prod_{p\geq 0}}[(p+1)n+pn^{\prime})]_{c/2,(p+1)b+pb^{\prime}}\biggr{)}\\ &\quad\times\biggl{(}\,\prod_{p\geq 0}[2^{p}(n+n^{\prime})]_{2^{p-1}c,2^{p}(b+b^{\prime})-2^{p-1}c}[2^{p}(n+n^{\prime})]_{2^{p-1}c,2^{p}(b+b^{\prime})+2^{p-1}c}\biggr{)}\\ &\quad\times\biggl{(}\,\buildrel\longleftarrow\over{\prod_{p\geq 0}}[pn+(p+1)n^{\prime})]_{c/2,pb+(p+1)b^{\prime}}\biggr{)}.\end{split}$
In particular, if we set $n=\mathbf{e}_{1}$, $n^{\prime}=\mathbf{e}_{2}$, $c=1$, and $b=b^{\prime}=0$, we obtain the relation (3.75).
Proof.
It is enough to prove that, for a given positive integer $k$, the relation holds modulo $G^{>\ell}$ with $\ell=k\deg(n+n^{\prime})-1$. We prove it by the induction on $k$. For $k=1$, the RHS reduced to
(A.8) $[n]_{c/2,b}[n^{\prime}]_{c/2,b^{\prime}}\mod G^{>\ell}.$
Then, the relation certainly holds because $[n]_{c/2,b}=[n]_{c,b-c/2}[n]_{c,b+c/2}$ and $[n^{\prime}]_{c/2,b^{\prime}}=[n^{\prime}]_{c,b^{\prime}-c/2}[n^{\prime}]_{c,b^{\prime}+c/2}$ commute modulo $G^{>\ell}$ by the pentagon relation. Suppose that the claim holds for $k-1$ for some $k\geq 2$. First, we consider the relation
(A.9) $\begin{split}&\ [n^{\prime}]_{c/2,b^{\prime}}[n]_{c/2,b}\\ =&\ ([n^{\prime}]_{c/2,b^{\prime}}[n]_{c,b-c/2})[n]_{c,b+c/2}\\ =&\ [n]_{c,b-c/2}[n+n^{\prime}]_{c/2,b+b^{\prime}-c/2}[n+2n^{\prime}]_{c,b+2b^{\prime}-c/2}([n^{\prime}]_{c/2,b^{\prime}}[n]_{c,b+c/2})\\ =&\ [n]_{c,b-c/2}[n+n^{\prime}]_{c/2,b+b^{\prime}-c/2}[n+2n^{\prime}]_{c,b+2b^{\prime}-c/2}\\ &\quad\times[n]_{c,b+c/2}[n+n^{\prime}]_{c/2,b+b^{\prime}+c/2}[n+2n^{\prime}]_{c,b+2b^{\prime}+c/2}[n^{\prime}]_{c/2,b^{\prime}},\end{split}$
where we used (A.1) twice for the parentheses. This calculation is essentially the same as (3.38). We need to calculate the anti-ordered product $[n+2n^{\prime}]_{c,b+2b^{\prime}-c/2}[n]_{c,b+c/2}$ in the last expression. Note that $\{n+2n^{\prime},n\}=2c$. Then, by the induction hypothesis, we have, modulo $G^{\ell^{\prime}}$ with $\ell^{\prime}=(k-1)\deg(n+(n+2n^{\prime}))$,
(A.10) $\begin{split}&\ [n+2n^{\prime}]_{c,b+2b^{\prime}-c/2}[n]_{c,b+c/2}\\ \equiv&\ \biggl{(}\,\buildrel\longrightarrow\over{\prod_{p\geq 0}}[(2p+1)n+2pn^{\prime})]_{c,(2p+1)b+pb^{\prime}+c/2}\biggr{)}\\ &\quad\times\biggl{(}\,\prod_{p\geq 0}[2^{p+1}(n+n^{\prime})]_{2^{p}c,2^{p+1}(b+b^{\prime})-2^{p}c}[2^{p+1}(n+n^{\prime})]_{2^{p}c,2^{p+1}(b+b^{\prime})+2^{p}c}\biggr{)}\\ &\quad\times\biggl{(}\,\buildrel\longleftarrow\over{\prod_{p\geq 0}}[(2p+1)n+(2p+2)n^{\prime}]_{c,(2p+1)b+(2p+2)b^{\prime}-c/2}\biggr{)}.\end{split}$
Meanwhile, we have the inequality
(A.11) $(k-1)\deg(n+(n+2n^{\prime}))=2(k-1)\deg(n+n^{\prime})\geq k\deg(n+n^{\prime}),$
where the second inequality holds due to the assumption $k\geq 2$. Therefore, the relation holds also modulo $G^{>\ell}$ with $\ell=k\deg(n+n^{\prime})$. We put (A.10) into the last expression of (A.9) and apply Lemma A.3 to terms therein as
(A.12) $\begin{split}&\ [n+n^{\prime}]_{c/2,b+b^{\prime}-c/2}\biggl{(}\,\buildrel\longrightarrow\over{\prod_{p\geq 0}}[(2p+1)n+2pn^{\prime}]_{c,(2p+1)b+2pb^{\prime}+c/2}\biggr{)}\\ =&\ [n]_{c,b+c/2}\biggl{(}\,\buildrel\longrightarrow\over{\prod_{p\geq 1}}[(p+1)n+pn^{\prime}]_{c/2,(p+1)b+pb^{\prime}}\biggr{)}[n+n^{\prime}]_{c/2,b+b^{\prime}-c/2},\end{split}$
(A.13) $\begin{split}&\ \biggl{(}\,\buildrel\longleftarrow\over{\prod_{p\geq 0}}[(2p+1)n+(2p+2)n^{\prime}]_{c,(2p+1)b+(2p+2)b^{\prime}-c/2}\biggr{)}[n+n^{\prime}]_{c/2,b+b^{\prime}+c/2}\\ =&\ [n+n^{\prime}]_{c/2,b+b^{\prime}+c/2}\biggl{(}\,\buildrel\longleftarrow\over{\prod_{p\geq 1}}[(p+1)n+(p+2)n^{\prime}]_{c/2,(p+1)b+(p+2)b^{\prime}}\biggr{)}[n+2n^{\prime}]_{c,b+2b^{\prime}-c/2}.\end{split}$
Then, after a little manipulation, we obtain the desired expression in the RHS of (A.1). ∎
A.2. Derivation of relation (3.103)
Lemma A.5 (cf. [Mat21, Lemma 4]).
Let $\ell$ be any nonnegative integer. If $\{n^{\prime},n\}=c>0$, the following formula holds:
(A.14) $\begin{split}&\ [n^{\prime}]_{c,b^{\prime}}\biggl{(}\,\buildrel\longrightarrow\over{\prod_{0\leq p\leq\ell}}[n+pn^{\prime}]_{c/2,b+pb^{\prime}\pm pc/2}\biggr{)}\\ =&\ [n]_{c/2,b}\biggl{(}\,\buildrel\longrightarrow\over{\prod_{1\leq p\leq\ell}}[2n+(2p-1)n^{\prime}]_{c,2b+(2p-1)b^{\prime}\pm(p-1)c}[n+pn^{\prime}]_{c/4,b+pb^{\prime}\pm(2p-1)/4c}\biggr{)}\\ &\quad\times[2n+(2\ell+1)n^{\prime}]_{c,2b+(2\ell+1)b^{\prime}\pm\ell c}[n+(\ell+1)n^{\prime}]_{c/2,b+(\ell+1)b^{\prime}\pm\ell c/2}[n^{\prime}]_{c,b^{\prime}},\end{split}$
(A.15) $\begin{split}&\ \biggl{(}\,\buildrel\longleftarrow\over{\prod_{0\leq p\leq\ell}}[pn+n^{\prime}]_{c/2,pb+b^{\prime}\pm pc/2}\biggr{)}[n]_{c,b}\\ =&\ [n]_{c,b}[(\ell+1)n+n^{\prime}]_{c/2,(\ell+1)b+b^{\prime}\pm\ell c/2}[(2\ell+1)n+2n^{\prime}]_{c,(2\ell+1)b+2b^{\prime}\pm\ell c}\\ &\quad\times\biggl{(}\,\buildrel\longleftarrow\over{\prod_{1\leq p\leq\ell}}[pn+n^{\prime}]_{c/4,pb+b^{\prime}\pm(2p-1)c/4}[(2p-1)n+2n^{\prime}]_{c,(2p-1)b+2b^{\prime}\pm(p-1)c}\biggr{)}[n^{\prime}]_{c/2,b^{\prime}}.\end{split}$
Proof.
For $\ell=0$, they coincide with the relations in Lemma A.1. Then, they are proved by the induction on $\ell$ with the pentagon relation and Lemma A.1 in the same way as [Mat21, Lemma 4]. ∎
Lemma A.6 (cf. [Mat21, Lemma 3]).
Let $\ell$ be any nonnegative integer. If $\{n^{\prime},n\}=c>0$, the following formulas hold:
(A.16) $\begin{split}&\ [n^{\prime}]_{c,b^{\prime}}\biggl{(}\,\buildrel\longrightarrow\over{\prod_{p\geq 0}}[n+pn^{\prime}]_{c/2,b+pb^{\prime}\pm pc/2}\biggr{)}\\ =&\ [n]_{c/2,b}\biggl{(}\,\buildrel\longrightarrow\over{\prod_{p\geq 1}}[2n+(2p-1)n^{\prime}]_{c,2b+(2p-1)b^{\prime}\pm(p-1)c}[n+pn^{\prime}]_{c/4,b+pb^{\prime}\pm(2p-1)/4c}\biggr{)}[n^{\prime}]_{c,b^{\prime}},\end{split}$
(A.17) $\begin{split}&\ \biggl{(}\,\buildrel\longleftarrow\over{\prod_{p\geq 0}}[pn+n^{\prime}]_{c/2,pb+b^{\prime}\pm pc/2}\biggr{)}[n]_{c,b}\\ =&\ [n]_{c,b}\biggl{(}\,\buildrel\longleftarrow\over{\prod_{p\geq 1}}[pn+n^{\prime}]_{c/4,pb+b^{\prime}\pm(2p-1)c/4}[(2p-1)n+2n^{\prime}]_{c,(2p-1)b+2b^{\prime}\pm(p-1)c}\biggr{)}[n^{\prime}]_{c/2,b^{\prime}}.\end{split}$
Proof.
This is obtained by taking the limit $\ell\rightarrow\infty$ of the relations in Lemma A.5. ∎
Theorem A.7 (cf. [Mat21, Theorem 3]).
If $\{n^{\prime},n\}=c>0$, the following formula holds:
(A.18) $\begin{split}&\ [n^{\prime}]_{c/4,b^{\prime}}[n]_{c,b}\\ =&\ \biggl{(}\,\buildrel\longrightarrow\over{\prod_{p\geq 0}}[(2p+1)n+4pn^{\prime}]_{c,(2p+1)b+4pb^{\prime}}[(p+1)n+(2p+1)n^{\prime}]_{c/4,(p+1)b+(2p+1)b^{\prime}}\biggr{)}\\ &\quad\times[n+2n^{\prime}]_{c/2,b+2b^{\prime}}\\ &\quad\times\biggl{(}\,\prod_{p\geq 0}[2^{p}(n+2n^{\prime})]_{2^{p-1}c,2^{p}(b+2b^{\prime})-2^{p-1}c}[2^{p}(n+2n^{\prime})]_{2^{p-1}c,2^{p}(b+2b^{\prime})+2^{p-1}c}\biggr{)}\\ &\quad\times\biggl{(}\,\buildrel\longleftarrow\over{\prod_{p\geq 0}}[(2p+1)n+(4p+4)n^{\prime}]_{c,(2p+1)b+(4p+4)b^{\prime}}[pn+(2p+1)n^{\prime}]_{c/4,pb+(2p+1)b^{\prime}}\biggr{)}.\end{split}$
In particular, if we set $n=\mathbf{e}_{1}$, $n^{\prime}=\mathbf{e}_{2}$, $c=1$, and $b=b^{\prime}=0$, we obtain the relation (3.75).
Proof.
In constrast to the proof of Theorem A.4, we do not use the induction; rather, we use Theorem A.4. First, we consider the relation
(A.19) $\begin{split}&\ [n^{\prime}]_{c/4,b^{\prime}}[n]_{c,b}\\ =&\ [n^{\prime}]_{c/2,b^{\prime}-c/4}([n^{\prime}]_{c/2,b^{\prime}+c/4}[n]_{c,b})\\ =&\ ([n^{\prime}]_{c/2,b^{\prime}-c/4}[n]_{c,b})[n+n^{\prime}]_{c/2,b+b^{\prime}+c/4}[n+2n^{\prime}]_{c,b+2b^{\prime}+c/2}[n^{\prime}]_{c/2,b^{\prime}+c/4}\\ =&\ [n]_{c,b}[n+n^{\prime}]_{c/2,b+b^{\prime}-c/4}[n+2n^{\prime}]_{c,b+2b^{\prime}-c/2}\\ &\quad\times[n^{\prime}]_{c/2,b^{\prime}-c/4}[n+n^{\prime}]_{c/2,b+b^{\prime}+c/4}[n+2n^{\prime}]_{c,b+2b^{\prime}+c/2}[n^{\prime}]_{c/2,b^{\prime}+c/4},\end{split}$
where we used (A.1) twice for the parentheses. We need to calculate the anti-ordered product $[n^{\prime}]_{c/2,b^{\prime}-c/4}[n+n^{\prime}]_{c/2,b+b^{\prime}+c/4}$ in the last expression. By Theorem A.4, we have
(A.20) $\begin{split}&\ [n^{\prime}]_{c/2,b^{\prime}-c/4}[n+n^{\prime}]_{c/2,b+b^{\prime}+c/4}\\ =&\ \biggl{(}\,\buildrel\longrightarrow\over{\prod_{p\geq 0}}[(p+1)n+(2p+1)n^{\prime}]_{c,(p+1)b+(2p+1)b^{\prime}+c/4}\biggr{)}\\ &\quad\times\prod_{p\geq 0}[2^{p}(n+2n^{\prime})]_{2^{p-1}c,2^{p}(b+2b^{\prime})-2^{p-1}c}[2^{p}(n+2n^{\prime})]_{2^{p-1}c,2^{p}(b+2b^{\prime})+2^{p-1}c}\\ &\quad\times\biggl{(}\,\buildrel\longleftarrow\over{\prod_{p\geq 0}}[pn+(2p+1)n^{\prime}]_{c,pb+(2p+1)b^{\prime}-c/4}\biggr{)}.\end{split}$
We put (A.20) into the last expression of (A.19) and apply Lemma A.5 to terms therein as
(A.21) $\begin{split}&\ [n+2n^{\prime}]_{c,b+2b^{\prime}-c/2}\biggl{(}\,\buildrel\longrightarrow\over{\prod_{p\geq 0}}[(p+1)n+(2p+1)n^{\prime}]_{c,(p+1)b+(2p+1)b^{\prime}+c/4}\biggr{)}\\ =&\ [n+n^{\prime}]_{c/2,b+b^{\prime}+c/4}\\ &\quad\times\biggl{(}\,\buildrel\longrightarrow\over{\prod_{p\geq 1}}[(2p+1)n+4pn^{\prime}]_{c,(2p+1)b+4pb^{\prime}}[(p+1)n+(2p+1)n^{\prime}]_{c/4,(p+1)b+(2p+1)b^{\prime}}\biggr{)}\\ &\quad\times[n+2n^{\prime}]_{c,b+2b^{\prime}-c/2},\end{split}$
(A.22) $\begin{split}&\ \biggl{(}\,\buildrel\longleftarrow\over{\prod_{p\geq 0}}[pn+(2p+1)n^{\prime}]_{c,pb+(2p+1)b^{\prime}-c/4}\biggr{)}[n+2n^{\prime}]_{c,b+2b^{\prime}+c/2}\\ =&\ [n+2n^{\prime}]_{c,b+2b^{\prime}+c/2}\\ &\quad\times\biggl{(}\,\buildrel\longleftarrow\over{\prod_{p\geq 1}}[pn+(2p+1)n^{\prime}]_{c/4,pb+(2p+1)b^{\prime}}[(2p-1)n+4pn^{\prime}]_{c,(2p-1)b+4pb^{\prime}}\biggr{)}\\ &\quad\times[n^{\prime}]_{c/2,b^{\prime}-c/4}.\end{split}$
Then, after a little manipulation, we obtain the desired expression in the RHS of (A.18). ∎
Appendix B Positivity at degree 4
This is a continuation of Example 5.9 for degree 4. The purpose of writing down the explicit expressions is just to demonstrate how the problem becomes complicated even for such a small degree. We have
(B.1) $\delta((3,1))$ $=k\delta_{1},$
(B.2) $\delta((2,2))$ =$=\begin{cases}k\delta_{1}/2&(k\delta_{1}:\textrm{even}),\\ k\delta_{1}&(k\delta_{1}:\textrm{odd}),\end{cases}$
(B.3) $\delta((1,3))$ =$=\begin{cases}k\delta_{1}/3&(k\equiv 0\mod 3),\\ k\delta_{1}&(k\not\equiv 0\mod 3).\end{cases}$
Firstly, we have the term for $n=(3,1)$,
(B.4) $\begin{split}\prod_{j=1}^{k\delta_{1}}\prod_{i=1}^{\delta_{1}-2}\prod_{t=1}^{i}\prod_{s=1}^{t}\begin{bmatrix}3\\ 1\end{bmatrix}_{1,\frac{-2k\delta_{1}+3k+1-2j+2ki+2kt+2ks}{k\delta_{1}}}\\ \qquad=\prod_{i=1}^{\delta_{1}-2}\prod_{t=1}^{i}\prod_{s=1}^{t}\begin{bmatrix}3\\ 1\end{bmatrix}_{\frac{1}{k\delta_{1}},\frac{-3\delta_{1}+3+2i+2t+2s}{\delta_{1}}}.\end{split}$
Next, we have the term for $n=(2,2)$,
(B.9) $\prod_{j=1}^{k\delta_{1}}\prod_{s=1}^{j-1}\prod_{i=1}^{\delta_{1}-1}\prod_{t=1}^{i}\begin{bmatrix}2\\ 2\end{bmatrix}_{1,\frac{-2j-2s+2ki+2kt+2}{k\delta_{1}}-1}\begin{bmatrix}2\\ 2\end{bmatrix}_{1,\frac{-2j-2s+2ki+2kt+2}{k\delta_{1}}+1},$
which is written as
(B.10) $\begin{split}&\prod_{s=1}^{k\delta_{1}-1}\prod_{i=1}^{\delta_{1}-1}\prod_{t=1}^{i}\begin{bmatrix}2\\ 2\end{bmatrix}_{\frac{2}{k\delta_{1}},\frac{-3k\delta_{1}+2s+2ki+2kt}{k\delta_{1}}-1}\begin{bmatrix}2\\ 2\end{bmatrix}_{\frac{2}{k\delta_{1}},\frac{-3k\delta_{1}+2s+2ki+2kt}{k\delta_{1}}+1}\\ &\hskip 220.0pt(k\delta_{1}:\text{even}),\end{split}$
(B.11) $\begin{split}&\prod_{s=1}^{(k\delta_{1}-1)/2}\prod_{i=1}^{\delta_{1}-1}\prod_{t=1}^{i}\begin{bmatrix}2\\ 2\end{bmatrix}_{\frac{1}{k\delta_{1}},\frac{-3k\delta_{1}-1+4s+2ki+2kt}{k\delta_{1}}-1}\begin{bmatrix}2\\ 2\end{bmatrix}_{\frac{1}{k\delta_{1}},\frac{-3k\delta_{1}-1+4s+2ki+2kt}{k\delta_{1}}+1}\\ &\hskip 250.0pt(k\delta_{1}:\text{odd}).\end{split}$
Here we used the following resummation formula for any positive integer $p$:
(B.12) $\begin{split}\sum_{j=1}^{p}\sum_{s=1}^{j-1}x^{-2j-2s}&=\frac{x^{-4p+2}(1-x^{2p-2})(1-x^{2p})}{(1-x^{2})(1-x^{4})}\\ &=\begin{cases}\displaystyle\sum_{j=1}^{p/2}\sum_{s=1}^{p-1}x^{-2p-4j+2s}&(p:\text{even}),\\ \displaystyle\sum_{j=1}^{p}\sum_{s=1}^{(p-1)/2}x^{-2p-2-2j+4s}&(p:\text{odd}).\end{cases}\end{split}$
Finally, we have the term for $n=(1,3)$,
(B.13) $\begin{split}\prod_{j=1}^{k\delta_{1}-2}\prod_{i=1}^{\delta_{1}}\prod_{t=1}^{k\delta_{1}-j-1}\prod_{s=1}^{t}\begin{bmatrix}1\\ 3\end{bmatrix}_{1,\frac{k+1-2j-2ki+2t+2s}{k\delta_{1}}},\end{split}$
which is written as
(B.16) $\prod_{i=1}^{\delta_{1}}\prod_{t=1}^{k\delta_{1}-2}\prod_{s=1}^{t}\begin{bmatrix}1\\ 3\end{bmatrix}_{\frac{3}{k\delta_{1}},\frac{-k\delta_{1}+k+2-2ki+2t+2s}{k\delta_{1}}}\quad(k\delta_{1}\equiv 0\mod 3),$
(B.17) $\begin{split}&\prod_{i=1}^{\delta_{1}}\prod_{t=1}^{(k\delta_{1}-1)/6}\Biggl{(}\begin{bmatrix}1\\ 3\end{bmatrix}_{\frac{1}{k\delta_{1}},\frac{-k\delta_{1}+k-4-2ki+12t}{k\delta_{1}}}\begin{bmatrix}1\\ 3\end{bmatrix}_{\frac{1}{k\delta_{1}},\frac{k\delta_{1}+k-6-2ki+12t}{k\delta_{1}}}\\ &\qquad\times\prod_{s=1}^{k\delta_{1}-4}\begin{bmatrix}1\\ 3\end{bmatrix}_{\frac{1}{k\delta_{1}},\frac{-k\delta_{1}+k-2-2ki+12t+2s}{k\delta_{1}}}\Biggr{)}\quad(k\delta_{1}\equiv 1\mod 6),\end{split}$
(B.18) $\begin{split}&\prod_{i=1}^{\delta_{1}}\prod_{t=1}^{(k\delta_{1}-2)/6}\Biggl{(}\begin{bmatrix}1\\ 3\end{bmatrix}_{\frac{1}{k\delta_{1}},\frac{-k\delta_{1}+k-4-2ki+12t}{k\delta_{1}}}\begin{bmatrix}1\\ 3\end{bmatrix}_{\frac{1}{k\delta_{1}},\frac{k\delta_{1}+k-4-2ki+12t}{k\delta_{1}}}\\ &\qquad\times\prod_{s=1}^{k\delta_{1}-3}\begin{bmatrix}1\\ 3\end{bmatrix}_{\frac{1}{k\delta_{1}},\frac{-k\delta_{1}+k-2-2ki+12t+2s}{k\delta_{1}}}\Biggr{)}\quad(k\delta_{1}\equiv 2\mod 6),\end{split}$
(B.21) $\prod_{i=1}^{\delta_{1}}\prod_{t=1}^{(k\delta_{1}-2)/2}\prod_{s=1}^{(k\delta_{2}-1)/3}\begin{bmatrix}1\\ 3\end{bmatrix}_{\frac{1}{k\delta_{1}},\frac{-k\delta_{1}+k-2-2ki+4t+6s}{k\delta_{1}}}\quad(k\delta_{1}\equiv 4\mod 6),$
(B.24) $\prod_{i=1}^{\delta_{1}}\prod_{t=1}^{(k\delta_{1}-2)/3}\prod_{s=1}^{(k\delta_{2}-1)/2}\begin{bmatrix}1\\ 3\end{bmatrix}_{\frac{1}{k\delta_{1}},\frac{-k\delta_{1}+k-2-2ki+6t+4s}{k\delta_{1}}}\quad(k\delta_{1}\equiv 5\mod 6).$
Here we used the following resummation formula for any positive integer $p$: The Laurent polynomial
(B.25) $\sum_{j=1}^{p-2}\sum_{t=1}^{p-j-1}\sum_{s=1}^{t}x^{-2j+2t+2s}=\frac{x^{-2p+8}(1-x^{2p-4})(1-x^{2p-2})(1-x^{2p})}{(1-x^{2})(1-x^{4})(1-x^{6})}$
equals to
(B.26) $\displaystyle\sum_{j=1}^{p/3}\sum_{t=1}^{p-2}\sum_{s=1}^{t}x^{4-6j+2t+2s}\quad(p\equiv 0\mod 3),$
(B.27) $\displaystyle\sum_{j=1}^{p}\sum_{t=1}^{(p-1)/6}x^{-2p-6+2j+12t}\biggl{(}1+x^{2p-2}+\sum_{s=1}^{p-4}x^{2+2s}\biggr{)}\quad(p\equiv 1\mod 6),$
(B.28) $\displaystyle\sum_{j=1}^{p}\sum_{t=1}^{(p-2)/6}x^{-2p-6+2j+12t}\biggl{(}1+x^{2p}+\sum_{s=1}^{p-3}x^{2+2s}\biggr{)}\quad(p\equiv 2\mod 6),$
(B.29) $\displaystyle\sum_{j=1}^{p}\sum_{t=1}^{(p-2)/2}\sum_{s=1}^{(p-1)/3}x^{-2p-4+2j+4t+6s}\quad(p\equiv 4\mod 6),$
(B.30) $\displaystyle\sum_{j=1}^{p}\sum_{t=1}^{(p-2)/3}\sum_{s=1}^{(p-1)/2}x^{-2p-4+2j+6t+4s}\quad(p\equiv 5\mod 6).$
Again, we need to resolve the discrepancy between (B.3) and (B.16) for $n=(1,3)$ when $\delta_{1}\equiv 0$ and $k\not\equiv 0$ mod 3, e.g., $(\delta_{1},\delta_{2})=(3,6)$. The positivity formula that is similar to (5.28) exists. However, it becomes much more complicated even for just writing it down. First, we note that, due to the translation invariance of the expression (B.16) with respect to the index $i$ therein, it is enough to concentrate on the case $\delta_{1}=3$. Let
(B.31) $\begin{split}A_{k}(x):=&\sum_{i=1}^{3}\sum_{t=1}^{3k-2}\sum_{s=1}^{t}x^{-2ki+2t+2s+6k-4}\\ =&\ \frac{(1-x^{6k-4})(1-x^{6k-2})(1-x^{6k})}{(1-x^{2})(1-x^{4})(1-x^{2k})}\end{split}$
be the (normalized) generating polynomial of the quantum data of the product (B.16) for $\delta_{1}=3$. Let
(B.32) $B_{k}(x):=\frac{A_{k}(x)}{1+x^{2}+x^{4}}=\frac{(1-x^{6k-4})(1-x^{6k-2})(1-x^{6k})}{(1-x^{4})(1-x^{6})(1-x^{2k})}.$
Then, the positivity property of the product (B.16) is equivalent to the fact that $B_{k}(x)$ is a polynomial in $x$ with nonnegative integer coefficients for $k\not\equiv 0$ mod 3. This is not obvious from the expression (B.32). However, one can find and prove such a positive expression with the help of a computer algebra software. (We used SageMath 9.4.) We separate it into the cases $k\equiv 1$, 2, 4, 5 mod 6.
(i). Let $k=6p+1$ ($p\in\mathbb{Z}_{\geq 0}$). Then, the polynomial $B_{k}(x)$ has the following positive expression.
$B_{6p+1}(x)$ $=\sum_{t=1}^{p}\sum_{s=1}^{3}t(x^{12t+4s-16}+x^{-12t-4s+16+96p})$
$\quad+\sum_{t=1}^{p}\bigl{\{}\sum_{s=1}^{2}(p+2t-1)(x^{12t+4s-16+12p}+x^{-12t-4s+16+84p})$
$\qquad\qquad+(p+2t)(x^{12t-4+12p}+x^{-12t+4+84p})\bigr{\}}$
$\quad+\sum_{t=1}^{3p}(3p+t)(x^{4t-4+24p}+x^{-4t+4+72p})$
$\quad+\sum_{t=1}^{p}\sum_{s=1}^{3}(6p+t)(x^{12t+4s-16+36p}+x^{-12t-4s+16+60p})$
$\quad+(7p+1)x^{48p}$
$\quad+\sum_{t=1}^{p-1}\sum_{s=1}^{3}t(x^{12t+4s-10}+x^{-12t-4s+10+96p})$
$\quad+\sum_{t=1}^{p+1}\bigl{\{}\sum_{s=1}^{2}(p+2t-2)(x^{12t+4s-22+12p}+x^{-12t-4s+22+84p})$
$\qquad\qquad+(p+2t-1)(x^{12t-10+12p}+x^{-12t+10+84p})\bigr{\}}$
$\quad+\sum_{t=1}^{3p-2}(3p+t+1)(x^{4t+2+24p}+x^{-4t-2+72p})$
$\quad+6p(x^{36p-2}+x^{36p+2}+x^{60p-2}+x^{60p+2})$
$\quad+\sum_{t=1}^{p-1}\sum_{s=1}^{3}(6p+t)(x^{12t+4s-10+36p}+x^{-12t-4s+10+60p})$
$\quad+7p(x^{48p-6}+x^{48p-2}+x^{48p+2}+x^{48p+6}).$
(ii). Let $k=6p+2$ ($p\in\mathbb{Z}_{\geq 0}$). Then, the polynomial $B_{k}(x)$ has the following positive expression.
$B_{6p+2}(x)$ $=\sum_{t=1}^{p}\sum_{s=1}^{3}t(x^{12t+4s-16}+x^{-12t-4s+32+96p})$
$\quad+\sum_{t=1}^{p}\bigl{\{}\sum_{s=1}^{2}(p+2t)(x^{12t+4s-12+12p}+x^{-12t-4s+28+84p})$
$\qquad\qquad+(p+2t-1)(x^{12t-12+12p}+x^{-12t+28+84p})\bigr{\}}$
$\quad+\sum_{t=1}^{3p+1}(3p+t)(x^{4t-4+24p}+x^{-4t+20+72p})$
$\quad+(6p+2)(x^{36p+4}+x^{36p+8}+x^{60p+8}+x^{60p+12})$
$\quad+\sum_{t=1}^{p-1}\sum_{s=1}^{3}(6p+t+2)(x^{12t+4s-4+36p}+x^{-12t-4s+20+60p})$
$\quad+(7p+2)(x^{48p}+x^{48p+4}+x^{48p+8}+x^{48p+12}+x^{48p+16})$
$\quad+\sum_{t=1}^{p}\sum_{s=1}^{3}t(x^{12t+4s-10}+x^{-12t-4s+26+96p})$
$\quad+\sum_{t=1}^{p}\bigl{\{}\sum_{s=1}^{2}(p+2t)(x^{12t+4s-6+12p}+x^{-12t-4s+22+84p})$
$\qquad\qquad+(p+2t-1)(x^{12t-6+12p}+x^{-12t+22+84p})\bigr{\}}$
$\quad+\sum_{t=1}^{3p}(3p+t)(x^{4t+2+24p}+x^{-4t+14+72p})$
$\quad+\sum_{t=1}^{p}\sum_{s=1}^{3}(6p+t)(x^{12t+4s-10+36p}+x^{-12t-4s+26+60p})$
$\quad+(7p+1)(x^{48p+6}+x^{48p+10}).$
(iii). Let $k=6p+4$ ($p\in\mathbb{Z}_{\geq 0}$). Then, the polynomial $B_{k}(x)$ has the following positive expression.
$B_{6p+4}(x)$ $=\sum_{t=1}^{p}\sum_{s=1}^{3}t(x^{12t+4s-16}+x^{-12t-4s+64+96p})$
$\quad+\sum_{t=1}^{p+1}\bigl{\{}\sum_{s=1}^{2}(p+2t-1)(x^{12t+4s-16+12p}+x^{-12t-4s+64+84p})$
$\qquad\qquad+(p+2t)(x^{12t-4+12p}+x^{-12t+52+84p})\bigr{\}}$
$\quad+\sum_{t=1}^{3p+1}(3p+t+2)(x^{4t+8+24p}+x^{-4t+40+72p})$
$\quad+(6p+4)(x^{36p+16}+x^{36p+20}+x^{60p+28}+x^{60p+32})$
$\quad+\sum_{t=1}^{p}\sum_{s=1}^{3}(6p+t+4)(x^{12t+4s+8+36p}+x^{-12t-4s+40+60p})$
$\quad+(7p+5)x^{48p+24}$
$\quad+\sum_{t=1}^{p}\sum_{s=1}^{3}t(x^{12t+4s-10}+x^{-12t-4s+58+96p})$
$\quad+\sum_{t=1}^{p+1}\bigl{\{}\sum_{s=1}^{2}(p+2t-1)(x^{12t+4s-10+12p}+x^{-12t-4s+58+84p})$
$\qquad\qquad+(p+2t)(x^{12t+2+12p}+x^{-12t+46+84p})\bigr{\}}$
$\quad+\sum_{t=1}^{3p}(3p+t+2)(x^{4t+14+24p}+x^{-4t+34+72p})$
$\quad+\sum_{t=1}^{p}\sum_{s=1}^{3}(6p+t+2)(x^{12t+4s+2+36p}+x^{-12t-4s+46+60p})$
$\quad+(7p+3)(x^{48p+18}+x^{48p+22}+x^{48p+26}+x^{48p+30}).$
(iv). Let $k=6p+5$ ($p\in\mathbb{Z}_{\geq 0}$). Then, the polynomial $B_{k}(x)$ has the following positive expression.
$B_{6p+5}(x)$ $=\sum_{t=1}^{p+1}\sum_{s=1}^{3}t(x^{12t+4s-16}+x^{-12t-4s+80+96p})$
$\quad+\sum_{t=1}^{p}\bigl{\{}\sum_{s=1}^{2}(p+2t+1)(x^{12t+4s+12p}+x^{-12t-4s+64+84p})$
$\qquad\qquad+(p+2t)(x^{12t+12p}+x^{-12t+64+84p})\bigr{\}}$
$\quad+\sum_{t=1}^{3p+3}(3p+t+1)(x^{4t+8+24p}+x^{-4t+56+72p})$
$\quad+\sum_{t=1}^{p}\sum_{s=1}^{3}(6p+t+4)(x^{12t+4s+8+36p}+x^{-12t-4s+56+60p})$
$\quad+(7p+5)(x^{48p+24}+x^{48p+28}+x^{48p+32}+x^{48p+36}+x^{48p+40})$
$\quad+\sum_{t=1}^{p}\sum_{s=1}^{3}t(x^{12t+4s-10}+x^{-12t-4s+74+96p})$
$\quad+\sum_{t=1}^{p+1}\bigl{\{}\sum_{s=1}^{2}(p+2t)(x^{12t+4s-6+12p}+x^{-12t-4s+70+84p})$
$\qquad\qquad+(p+2t-1)(x^{12t-6+12p}+x^{-12t+70+84p})\bigr{\}}$
$\quad+\sum_{t=1}^{3p+2}(3p+t+2)(x^{4t+14+24p}+x^{-4t+50+72p})$
$\quad+(6p+4)(x^{36p+26}+x^{60p+38})$
$\quad+\sum_{t=1}^{p}\sum_{s=1}^{3}(6p+t+4)(x^{12t+4s+14+36p}+x^{-12t-4s+50+60p})$
$\quad+(7p+5)(x^{48p+30}+x^{48p+34}).$